### v2

# Cryptocurrency as an Investable Asset Class: Coming of Age

Abstract
Cryptocurrencies are coming of age. We organize empirical regularities into ten stylized facts and analyze cryptocurrency through the lens of empirical asset pricing. We find important similarities with traditional markets—risk-adjusted performance is broadly comparable, and the cross-section of returns can be summarized by a small set of factors. However, cryptocurrency also has its own distinct character: jumps are frequent and large, and blockchain information helps drive prices. This common set of facts provides evidence that cryptocurrency is emerging as an investable asset class.
1 INTRODUCTION
This survey establishes ten key empirical regularities of cryptocurrencies motivated by the view that cryptocurrency is an asset and thus can and should be analyzed from the asset-pricing point of view and, more specifically, from the cryptocurrency asset pricing view. Although the literature is relatively young—cryptocurrencies trace back to Nakamoto (2008)’s short paper on Bitcoin—it is now at the point of maturity where it is useful to establish a core of stylized facts in cryptocurrency markets that can be used for both empirical and theoretical work. This common set of facts also provides emerging evidence that cryptocurrency is becoming an investable asset class.
Money has long been a central topic in economics—for example, see classic contributions of Tobin (1965) and Friedman (1968) to more recent advances. A large literature studied why money exists and how it functions—the role of money in economies with spatially separated agents (Townsend 1980), the coexistence of different types of money under different institutional restrictions (Wallace 1983), search-based models (Kiyotaki & Wright, 1989; Trejos & Wright, 1995; Shi, 1995; Lagos & Wright, 2005), mechanism-design approaches (Green & Zhou, 1998), the “money-as-memory” view (Kocherlakota, 1998), commodity and token money (Sargent & Wallace, 1983; Sargent, 2019), competing currencies (Kareken & Wallace, 1981), and the money-like safe assets (Gorton & Pennacchi 1990; Holmström & Tirole 1998).
When the paper of Nakamoto (2008) introduced a permissionless blockchain that provides public record-keeping without a central authority, financed and governed by a native token, the canonical literature on money had reached a degree of maturity. Permissionless ledgers can be viewed as technologies that supply public memory, with token systems financing and governing that record—keeping, and providing histories, communication, and enforcement mechanisms that are required for credit. With the rise of cryptocurrencies, these new ideas began to emerge primarily from practice more than theory and reopened classic monetary questions—record-keeping, commitment, enforcement—but now viewed from the point of view of blockchain economics.
When practice advances faster than theory, it is natural to start asking empirical questions that can themselves later be used to shape theory. The paper of Liu & Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing. A sizable literature uses an asset pricing lens to study risk exposures, cross-sectional return patterns, cash-flow-like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020, Liu et al. 2022, Alvarez et al. 2023, Sockin & Xiong 2023a, von Luckner et al. 2023, Kogan et al. 2024).
Many have viewed cryptocurrencies—and, not unjustifiably, many still view them—as hype, speculation, or fraud. When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system. [^1] Foley et al. (2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations. It is worth noting that several important financial innovations, such as the limited-liability company and mortgage-backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018, Leshno & Strack 2020, Biais et al. 2019, Kogan et al. 2022, Brunnermeier et al. 2019, Benigno et al. 2022, Capponi et al. 2025, Jermann & Xiang 2025), regulation (Cong & He 2019,Townsend 2020, Auer et al. 2025), digital asset pricing and valuation (Athey et al. 2016, Cong et al. 2021b, Biais et al. 2023), transaction costs (Easley et al. 2019, Cong et al. 2021a, Basu et al. 2023), central bank money (Auer et al. 2022, Whited et al. 2023, Schilling et al. 2024), securitization (Cong et al. 2022, Sockin & Xiong 2023a, Rogoff & You 2023, Goldstein et al. 2024), safe assets (Ma et al. 2025), international capital flow (von Luckner et al. 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al. 2024).
We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of ten stylized empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes. These empirical facts are also useful to evaluate and revisit both the classic and the new theories of money.
Ten Facts in Cryptocurrency Asset Pricing and Empirical Finance. Facts 1 and 2 overview the core asset-pricing characteristics of cryptocurrency markets—the risk-return trade-off and the diversification benefits. Facts 3 and 4 discuss the structure of returns in cryptocurrency—factors, or “smart-beta”, which explain the cross-section of returns and optimal portfolio shares. Fact 5 shows the importance of large, discrete price movements—jumps—and the importance of extreme events for understanding cryptocurrency risk and valuation. Fact 6 shows that the higher-order terms of a parsimonious four-factor model account for a large fraction of the cross-section of returns and provide a “glass box” alternative to the “black box” neural net models with interpretable economic factor structure. Fact 7 shows the importance of the blockchain structure and network economy phenomena for cryptocurrency prices—on-chain user adoption influences cryptocurrency returns. Facts 8 documents large and persistent market inefficiencies due to frictions that limit arbitrage, and discuss risks for arbitrageurs. Fact 9 is about the cryptocurrency futures market and funding premia, and their evolution. Finally, Fact 10 is about the importance of regulation, disclosure, and transparency. In sum, these ten facts can be viewed as providing evidence that cryptocurrency, and more broadly, digital assets are coming of age as an emerging investable asset class.
2 DATA
The empirical analysis is based on the most recent data aggregated from multiple sources.
First, we collect comprehensive price data on all cryptocurrencies from CoinGecko.com, covering both active and inactive coins between December 31, 2013, and September 6, 2025. The sample includes 18,622 coins listed on CoinGecko at the end of the period, as well as 29,230 inactive coins that have since been delisted but for which historical data are available. Including inactive coins allows us to avoid survivorship bias. We exclude stablecoins because they are designed to track the value of a currency, a basket of fiat currencies, or another reference asset. We also exclude wrapped tokens (such as Wrapped Ether) as they are essentially replicas of existing coins on different blockchains. Furthermore, we drop daily observations with missing or non-positive values for price, market capitalization, or trading volume. We also exclude all coins with fewer than 30 daily observations. The final dataset comprises 16,468 unique coins and approximately 9.5 million records.
Second, for the construction of the crypto carry strategy of Schmeling et al. (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025. The data on new addresses are from Liu et al. (2021). Moreover, for the construction of the crypto arbitrage strategy of Borri & Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025. Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.222Finally, we obtain price indices for the U.S. stock market, U.S. corporate and government bonds, a broad commodity index, gold, the U.S. dollar index, the Dow Jones US Real Estate Index, the U.S. constant maturity 10-year nominal and real bond yields from Bloomberg. The corresponding Bloomberg tickers are SPX Index, LUACTRUU Index, LUATTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
Unless otherwise noted, we aggregate daily closing prices to construct weekly coin returns. Each year is divided into 52 weeks: the first week contains the year’s first seven days, the next 50 weeks consist of seven days each, and the final week contains the remaining days of the year. For the construction of weekly returns, we exclude coins with a market capitalization below \$1 million at the formation date and winsorize returns at the 0.0025% level in both tails. We construct a cryptocurrency market return as the value-weighted return of all underlying coins. The cryptocurrency excess market return is constructed as the difference between the cryptocurrency market return and the risk-free rate measured as the 1-month T-bill rate at the weekly frequency, which we obtain from the Federal Reserve Bank of St. Louis.
Figure 1 summarizes the time-series evolution of cryptocurrency prices and market capitalization. A one-dollar investment at the start of the sample grows to about \$160 in Bitcoin and \$80 in the overall market. There is substantial variation over time. Market capitalization reaches nearly \$4 trillion, with Bitcoin and Ethereum accounting for approximately 58% and 12% of the total at the end of the sample.
(a) Return
(b) Capitalization
Figure 1: Cryptocurrency market returns and capitalization
Panel A of the figure plots the cryptocurrency market return against Bitcoin. Panel B plots the cryptocurrency market capitalization against the top five coins by market capitalization at the end of the sample (Bitcoin, Ethereum, Ripple, Binance Coin, and Solana). For returns, the figure shows the value of investment over time for one dollar of investment starting in December 31, 2013. For market capitalization, the figure shows the value over time of the outstanding coins in US dollar billions. Data end on September 6, 2025.
3 THE TEN FACTS OF CRYPTOCURRENCY
We organize the main discussions of this review into the ten key stylized empirical facts on cryptocurrency markets.
3.1 Fact 1: High return, high volatility—normal Sharpe ratio
Fact 1: The cryptocurrency market has an order of magnitude higher returns and volatility than equities, but a broadly similar Sharpe ratio.
Popular discussion of cryptocurrencies tends to focus either on their dramatic average returns or on dramatic volatility. From the classic finance point of view, the relevant measure is, however, risk-adjusted performance—the Sharpe ratio, i.e., mean excess return per unit of volatility (Sharpe 1966).
Liu & Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable. Fact 1 revisits and confirms this finding in a longer and larger dataset up to September 2025.
Table 1, Panel A, reports summary statistics for the sample beginning in 2014 and ending in September 2025, as well as for the samples pre-2020 and post-2020. We compare returns on the aggregate coin market, a large-cap index of the top 100 cryptocurrencies (Coin 100), Bitcoin, and the aggregate U.S. stock market. The average returns and volatilities of all cryptocurrency return series are an order of magnitude greater than those of the stock market across all sample periods. For instance, average weekly returns exceed 1% for all cryptocurrency indices, compared with less than 0.3% for the aggregate stock market. However, the weekly Sharpe ratios are similar across cryptocurrencies and equities. This, confirms that the Sharpe ratio calculated by Liu & Tsyvinski (2021) holds in the most recent sample.
Table 1, Panel A, reports maximum drawdowns and run-ups for the four return indices. Drawdowns are important as they provide a measure of tail-risk not fully reflected by standard deviation. Drawdowns and run-ups are substantially larger for cryptocurrencies than for equities. Importantly, for all three cryptocurrency indices, maximum drawdowns and run-ups are smaller in magnitude in the post-2020 sample than in the pre-2020 sample, while the reverse holds for equities. This pattern can also be observed in the standard deviations of returns—they decline for cryptocurrencies and rise for equities in the more recent period.
Table 1: Summary statistics
Panel A. Summary statistics returns
Mean (%) SD (%) t-Stat Sharpe Skew Kurt Max DD (%) Max RU (%)
(1) (2) (3) (4) (5) (6) (7) (8)
Full sample
Coin market 1.23 10.27 2.95 0.12 0.15 1.89 -89.82 227.74
Coin 100 1.24 10.30 2.98 0.12 0.15 1.84 -89.23 224.08
Bitcoin 1.33 10.12 3.24 0.13 0.33 2.16 -83.62 208.94
Stock 0.27 2.18 3.01 0.11 -0.66 3.95 -27.22 17.55
Pre-2020 sample
Coin market 1.20 11.37 1.87 0.10 0.25 1.39 -89.82 227.74
Coin 100 1.20 11.44 1.86 0.10 0.23 1.29 -89.23 224.08
Bitcoin 1.35 11.28 2.12 0.12 0.31 1.46 -83.62 208.94
Stock 0.24 1.79 2.35 0.12 -0.68 2.54 -17.09 10.67
Post-2020 sample
Coin market 1.26 9.00 2.41 0.13 -0.07 2.34 -77.57 83.62
Coin 100 1.29 8.98 2.46 0.14 -0.03 2.41 -76.78 84.36
Bitcoin 1.30 8.75 2.56 0.14 0.32 3.03 -74.94 95.10
Stock 0.29 2.53 2.00 0.10 -0.63 3.47 -27.22 17.55
Panel B. Coin market return correlations
Sample Stock Corp Govt Comm Gold Dollar Inflation
(1) (2) (3) (4) (5) (6) (7)
Full 0.20 0.08 0.03 0.14 0.12 -0.06 0.17
(4.92) (2.04) (0.70) (3.60) (2.87) (-1.49) (4.35)
Pre 2020 0.02 0.04 0.02 0.03 0.06 -0.03 0.03
(0.44) (0.66) (0.42) (0.51) (1.13) (-0.48) (0.46)
Post 2020 0.37 0.13 0.04 0.27 0.18 -0.10 0.31
(6.91) (2.20) (0.60) (4.74) (3.10) (-1.80) (5.66)
Notes: This table summarizes the main properties of coin market return. Panel A reports weekly summary statistics of the coin market index and compares them with the value-weighted index of the top 100 coins by market capitalization (Coin 100), Bitcoin and U.S. stock market returns. Reported measures include the mean, standard deviation, $t$-statistic, Sharpe ratio, skewness, kurtosis, maximum drawdown and run-up. Panel B presents pairwise correlation coefficients between coin market returns and returns on the U.S. stock market, corporate bonds, government bonds, a commodity index, gold, the dollar index, and the changes in inflation expectations. $t$-statistics are reported in parentheses. The full sample covers the period from January 7, 2014 to September 6, 2025. The pre-2020 sample is the sample before Jan 1, 2020, and the post-2020 sample is the sample since Jan 1, 2020.
3.2 Fact 2: Cryptocurrency is a distinct asset: it moves largely on its own—but increasingly with equities, gold and commodities
Fact 2: The aggregate cryptocurrency market remains largely uncorrelated with most traditional asset classes, but its correlation with U.S. equities has increased sharply—from 2% before 2020 to 37% after 2020.
How cryptocurrencies co-move with traditional asset classes is a first-order question for asset pricing and portfolio choice (e.g., Markowitz 1952). Correlations are an important determinant of diversification benefits and show whether cryptocurrency risk is distinct to that market or simply represent familiar risks from other markets. We revisit the cryptocurrency aggregate correlation structure using updated data, contrasting the pre-2020 and post-2020 return correlation.
One of the main results of Liu & Tsyvinski (2021) is that aggregate cryptocurrency returns have insignificant exposures to the returns of traditional asset classes, such as equities, bonds, commodities, and foreign exchanges. In other words, the correlation between the aggregate cryptocurrency returns and those of traditional asset classes are low. Fact 2 revisits this finding. We show that there is a rising correlation with the U.S. stock market.
Table 1, Panel B, reports correlations of aggregate cryptocurrency returns with those of major asset classes. In the pre-2020 sample, correlations with equities, corporate bonds, government bonds, commodities, gold, and the dollar are all statistically insignificant, confirming the earlier findings of Liu & Tsyvinski (2021). In contrast, in the post-2020 sample, aggregate cryptocurrency returns become significantly correlated with equities (37%), commodities (27%), gold (18%), and corporate bonds (13%) at the 5% level. The correlation with the dollar index is negative (-10%) and significant at the 10% level. Government bonds remain the only asset class with no significant link.
The most striking change is the correlation with the U.S. stock market, which rises from only 2% before 2020 to 37% afterwards. This substantial increase is an important indicator that cryptocurrencies have become more integrated with equity markets. We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S. Treasury bonds with 10-year maturity. Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years. In our view, this link with expected inflation is important, as it provides empirical evidence on the original motivation of cryptocurrencies as assets designed to hedge against the inflationary risks of government printing money.
The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong et al. (2023) and Jiang et al. (2025). Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization. The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market.
3.3 Fact 3: Significant diversification benefits from adding small amount of cryptocurrencies in investment portfolios
Fact 3: Cryptocurrency idiosyncratic risk can be significantly reduced by constructing optimal mean-variance asset portfolios that include cryptocurrencies—the inclusion of cryptocurrencies enlarges the mean-variance frontier. Optimal portfolio allocates 3.1% cryptocurrency for retail investors and 5.5% for institutional investors.
Modern finance theory, starting from Markowitz (1952), emphasizes the central insight of diversification: an asset’s covariance with the investor’s portfolio matters more than its volatility. Low correlations push the efficient frontier outward, and volatile assets can add value if their expected returns compensate for marginal risk. Since Fact 1 establishes that cryptocurrencies have low correlations with stocks and bonds, they can be effective risk diversifiers under standard portfolio-theory frameworks.
Liu & Tsyvinski (2021) show that cryptocurrencies can improve portfolio efficiency because of their historically low correlations with traditional asset classes and comparatively high expected returns. Using the Black–Litterman portfolio allocation model (Black & Litterman 1990) to incorporate investor views, Liu & Tsyvinski (2021) estimate that a representative investor would optimally allocate between 1% and 6% of wealth to cryptocurrencies, depending on the strength of those views. A complementary view comes from Bayesian portfolio theory. Duchin et al. (2025) obtain similar optimal weights in the 1%–5% range. Ang et al. (2023) find much larger optimal weights when investors have a preference for positive skewness, such as under constant relative risk aversion.
Because the optimal crypto weight is sensitive to the joint return distribution—especially the correlation between cryptocurrencies and traditional assets—and because this correlation structure shifted after 2020 (see Fact 2), we re-estimate the theoretical optimal allocations using post-2020 data under the same Black–Litterman framework.
Moreover, we extend the analysis to two representative investor classes: a retail investor and an institutional investor. In calculating the theoretical optimal crypto portfolio shares for them, we differ only in their investable options. This simplification abstracts from practical complications, such as liquidity, custody, and regulatory constraints, to make the exercise tractable.
For a retail investor, we start with a baseline portfolio that is 50–50 between financial assets and real estate, with the financial portion 60–40 stocks–bonds. Based on the model, if we add an allocation to the cryptocurrency market, we would raise the portfolio’s Sharpe ratio from 0.39 to 0.52, with an optimal crypto weight of 3.1% of total wealth. As a historical counterfactual, we also compute the theoretical optimal portfolio using only pre-2020 data; the optimal crypto share is 1.4%. Mapping this to per-capita U.S. net wealth in 2019, a 1.4% allocation implies an investment of about \$3,500; held from 2020 to the end of our sample, this position would have grown by roughly \$52,000 (about 1,462%). This is a back-of-the-envelope, buy-and-hold calculation that assumes no rebalancing.
For an institutional investor, we consider a broader investable universe—equities, government and corporate bonds, commodities, and real estate—with baseline weights proportional to each asset class’s market capitalization. Solving for the optimal portfolio and assuming slightly greater confidence in views on cryptocurrencies than in the retail case, the model implies an increase in the portfolio’s Sharpe ratio from 0.69 to 0.95, with an optimal share of 5.5% in the cryptocurrency market.
These results show that—even with the post-2020 increase in comovement with equities—optimal allocations to cryptocurrency are sizable and in the 3–5% range, with larger weights for institutional portfolios. It is important to note that the estimations in this section are back-of-the-envelope and rely on the properties of returns that may not persist and abstract from important issues related to liquidity, custody, and regulatory risk, and therefore should be viewed as theoretical upper bounds on the potential diversification benefits of this emerging investable asset class, not as actual investment recommendations. For context, CalPERS, one of the largest U.S. pension funds, reports portfolio shares of comparable magnitude in emerging-market debt (5.3%), high-yield bonds (5.3%), private debt (3.8%), and mortgage-backed securities (5.8%) as of June 2025, indicating that the model-implied weights of cryptocurrencies are within a realistic range.
3.4 Fact 4: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
Fact 4: A small number of smart beta strategies—based on crypto-size, crypto-momentum, and crypto-value—span the cross-section of cryptocurrency returns.
A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors. In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French (1992))—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies. For cryptocurrencies, Liu et al. (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns. Liu et al. (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power. We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE). [^3] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi et al. 2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al. 2022b; Borri & Santucci de Magistris 2022). The return to each predictor is measured as the long-short spread between the highest and lowest quintiles (Portfolio 5 minus Portfolio 1).
Table 2, Panels A and B, report average excess returns to the quintile portfolios and the corresponding long-short strategies for the full sample and the post-2020 subsample. All three predictors generate statistically significant long-short returns in the full sample. The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu et al. (2022). Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio. High bracket is the low-value coins and low bracket is the high-value coins. The sorting is done in week $t$ and the returns are the value-weighted average in the following week.
Importantly, these results remain robust in the post-2020 sample, the core set of smart beta strategies in the C-3 factor model and C-4 factor model continues to explain variation in cryptocurrency returns even in the more recent period.
The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns. For example, Asness et al. (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities. The C-4 factor model of Liu et al. (2022) and Liu et al. (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities. It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets. CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities). In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu et al. (2021) and proxy CVALUE with the price-to-new-address ratio. [^4] Liu et al. (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong et al. 2021b, Sockin & Xiong 2023a, b).
A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”. Specifically, Liu et al. (2022) and Liu et al. (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng et al. 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model. In Table 3, we study long-short strategies based on these signals post–2020. These strategies include portfolios based on long-horizon momentum, volatility, volume, and beta. We find that these long-short strategies continue not to generate significant average returns in the recent sample. [^5] In a recent study, Fieberg et al. (2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market. They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not. These results give evidence that the cross-section of cryptocurrency returns can be summarized by a small number of factors.
Table 2: Cryptocurrency smart beta strategies
Panel A: Full sample
Quintiles
1 2 3 4 5 5–1
CSIZE Low High
Mean 0.035*** 0.015* 0.014* 0.007 0.012*** -0.023***
tt(Mean) (3.27) (1.88) (1.76) (0.99) (2.59) (-2.66)
CMOM Low High
Mean -0.002 0.005 0.001 0.018*** 0.024*** 0.026***
tt(Mean) (-0.28) (0.68) (0.21) (2.84) (3.19) (3.89)
CVALUE Low High
Mean 0.034*** 0.026** 0.018** 0.014 -0.000 -0.035***
tt(Mean) (2.82) (2.24) (2.00) (1.61) (-0.05) (-3.50)
Panel B: Post-2020 sample
Quintiles
1 2 3 4 5 5–1
CSIZE Low High
Mean 0.021** 0.011 0.012 0.013 0.012** -0.009*
tt(Mean) (2.53) (1.29) (1.33) (1.42) (2.11) (-1.69)
CMOM Low High
Mean 0.001 0.001 0.009 0.016** 0.021** 0.021***
tt(Mean) (0.08) (0.17) (1.38) (2.41) (2.50) (3.70)
CVALUE Low High
Mean 0.030** 0.021* 0.017** 0.012* 0.007 -0.023**
tt(Mean) (2.35) (1.95) (2.06) (1.79) (1.35) (-2.32)
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for crypto smart beta strategies. The smart beta predictors are market capitalization (CSIZE), past two-week return (CMOM), and price-to-new address ratio (CVALUE). The long-short portfolio is long portfolio 5 and short portfolio 1 for all predictors. The mean returns are the time-series averages of weekly value-weighted portfolio excess returns. The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.
Table 3: Cryptocurrency smart beta strategies that are not beta
Panel A: Full sample
Quintiles
1 2 3 4 5 5–1
MOM12 Low High
Mean 0.015** 0.011 0.011* 0.011* 0.018*** 0.003
tt(Mean) (2.11) (1.52) (1.73) (1.84) (2.83) (0.47)
MOM24 Low High
Mean 0.021*** 0.018** 0.013** 0.016** 0.011* -0.009
tt(Mean) (2.81) (2.37) (2.22) (2.48) (1.96) (-1.50)
VOL Low High
Mean 0.013*** 0.022*** 0.014** 0.015* 0.011 -0.003
tt(Mean) (2.93) (2.86) (2.27) (1.95) (1.36) (-0.45)
VOLUME Low High
Mean 0.043** 0.012* 0.011 0.013* 0.013*** -0.030
tt(Mean) (2.17) (1.66) (1.51) (1.84) (2.77) (-1.60)
CMKT BETA Low High
Mean 0.009 0.012** 0.013** 0.016** 0.016** 0.007
tt(Mean) (1.47) (2.17) (2.35) (2.50) (2.38) (1.34)
MKT BETA Low High
Mean 0.006 0.012** 0.015*** 0.018** 0.017* 0.011*
tt(Mean) (0.88) (2.19) (2.81) (2.48) (1.96) (1.80)
Panel B: Post-2020 sample
1 2 3 4 5 5–1
MOM12 Low High
Mean 0.009 0.009 0.011 0.011* 0.016** 0.006
tt(Mean) (1.18) (1.15) (1.59) (1.84) (2.24) (1.30)
MOM24 Low High
Mean 0.014* 0.011 0.011 0.013* 0.011* -0.004
tt(Mean) (1.72) (1.52) (1.59) (1.90) (1.75) (-0.66)
VOL Low High
Mean 0.012** 0.011 0.014* 0.011 0.005 -0.007
tt(Mean) (2.23) (1.61) (1.70) (1.26) (0.59) (-1.39)
VOLUME Low High
Mean 0.047* 0.004 0.017 0.012 0.012** -0.035
tt(Mean) (1.69) (0.67) (1.60) (1.63) (2.16) (-1.35)
CMKT BETA Low High
Mean 0.006 0.013** 0.014** 0.009 0.014* 0.008
tt(Mean) (1.43) (2.02) (2.06) (1.37) (1.86) (1.64)
MKT BETA Low High
Mean 0.004 0.012** 0.013** 0.011* 0.011 0.007
tt(Mean) (0.61) (2.00) (2.00) (1.67) (1.27) (1.17)
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for crypto smart beta strategies based on insignificant predictors. The predictors are past 12-week and 24-week return (MOM12 and MOM24), past 24-week return standard deviation (VOL), past 24-week dollar trading volume (VOLUME), coin market beta (CMKT BET) and stock market beta (MKT BETA) estimated using past 24-week returns. The long-short portfolio is long portfolio 5 and short portfolio 1 for all predictors. The mean returns are the time-series averages of weekly value-weighted portfolio excess returns. The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.
3.5 Fact 5: Mind the Jumps—large, sudden price moves and “common disasters”
Fact 5: Jumps are frequent and account for a large share of variation of cryptocurrency returns, far more than in equities. Thus, disasters are not rare but common.
Leibniz wrote that natura non facit saltus—nature makes no jumps. In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities. Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973), in their celebrated option pricing paper, to model prices as Brownian motions. [^6] Merton (1976) is aware of potential jumps in equity prices and include the feature in his paper. Rare disasters are also used as a central ingredient to explain risk-premia (e.g., Rietz (1988); Barro (2006)). Cryptocurrencies are different: jumps, and hence disasters, are not “rare disasters” but are “common disasters.” Specifically, Borri & Santucci de Magistris (2022) show that 12.8% of Bitcoin’s daily variability can be attributed to jumps—large, sudden price moves—and that higher-order moments such as skewness and kurtosis explain the risk premium on Bitcoin. [^7] See early evidence from Scaillet et al. (2020) similarly documenting very frequent jumps in Bitcoin. Fact 5 revisits and extend this finding, documenting the frequency and size of jumps in cryptocurrency and showing that those are far more relevant than for the U.S. stock market.
Table 4 reports the frequency and magnitude of jumps in Bitcoin and Ethereum—the two largest cryptocurrencies by market capitalization—compared with U.S. stocks. Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006), a jump-robust measure of integrated variance. The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency. For U.S. equities, the same procedure is applied, and values represent cross-sectional averages across the constituents of the Dow Jones Industrial Average. Only jumps that are statistically significant at the 0.1% confidence level according to the BNS test of Barndorff-Nielsen & Shephard (2006) are retained. For Bitcoin and Ethereum, the probability of a significant jump on a given day is around 40 percent in the pre–2020 sample and 15–20 percent in the post–2020 period. In sharp contrast, the probability of a daily jump in U.S. stocks is below 10 percent before 2020 and falls to about 3 percent thereafter. Thus, jumps are far more frequent in cryptocurrencies than in equities. Conditional on a jump occurring, however, the share of total return variation explained by the jump is broadly similar across asset classes, particularly in the post–2020 period. This shows that what sets cryptocurrencies apart is the high frequency of jumps—contrary to Leibniz’s dictum that natura non facit saltus, cryptocurrencies are markets where saltus sunt regula (jumps are the rule), rather than the exception.
These results show that higher-order moments are central to understanding cryptocurrency risk. Equity and bond markets follow relatively smooth dynamics with only rare discontinuities. Cryptocurrency markets are characterized by frequent and sizable jumps that are a defining feature of their return distribution. This prevalence of jumps, and the corresponding role of higher-order moments, means significant challenges for risk management and portfolio allocation (see, e.g., Borri 2019). The implication is that we need empirical asset-pricing strategies that explicitly accommodate higher-order moments and nonlinearities (see, e.g., Kozak et al. 2020, Chetverikov et al. Forthcoming, Borri et al. 2025).
Table 4: Jumps in Cryptocurrency
Pr(J) N(J) Quantiles of J/RV
q(50) q(75) q(90) q(95)
Full sample
Bitcoin 28.005 946 31.237 44.322 58.819 67.419
Ethereum 25.992 878 34.420 56.584 74.992 85.270
Stock 5.313 137 37.743 44.271 52.009 57.310
Pre-2020 sample
Bitcoin 40.041 585 36.410 51.457 64.727 71.562
Ethereum 39.699 580 47.651 66.449 80.682 90.073
Stock 5.244 66 37.872 43.985 51.658 55.976
Post-2020 sample
Bitcoin 18.841 361 26.161 33.197 42.350 48.892
Ethereum 15.501 297 24.545 30.736 39.429 45.667
Stock 5.643 60 37.959 44.838 52.434 57.079
This table reports the probability of observing a statistically significant jump (P(J)), the number of such events (N(J)), and quantiles of the ratio of jump variation to realized variance (J/RV), conditional on a significant jump, for Bitcoin, Ethereum, and U.S. stocks. Significant jumps are identified using the BNS test with a 0.1% significance level. For U.S. equities, values are cross-sectional averages across the constituents of the Dow Jones Industrial Average. Data are from January 1, 2015 to March 31, 2025. The pre-2020 sample is the sample ending Jan 1, 2020. The post-2020 sample is the sample since Jan 1, 2020.
3.6 Fact 6: Few factors, higher orders—rather than machine learning: why less is more
Fact 6: In cryptocurrency, higher-order terms of the C-4 model accounts for approximately 25% of the cross-section of returns and captures most of the insights of machine learning methods.
A growing literature in finance uses machine learning methods, such as neural networks, to investigate the cross-section of returns (see, for example, Gu et al. 2020). Neural networks have become extremely successful in many prediction problems and moreover have the universal approximation property (Cybenko 1989; Hornik et al. 1989)—a sufficiently large neural network can approximate any continuous function—and thus, in principle, capture any features of the data. These ideas have been present in the machine learning literature for several decades, but only recently have they become dramatically successful, as computational advances enabled the estimation of very large models, essentially aligning with the core idea of the universal approximation theorem. This has led finance and economics, as well as many other fields, to develop models with a very large number of parameters. The main shortcoming of neural networks for finance is that they are “black box” models. Once millions and sometimes hundred of millions of “factors” are run through often rather complicated and delicate gradient descent-like algorithms, it is difficult to see what features are economically important, how exactly the model produces results, and if an environment changes, whether the model still continues to perform. [^8] For a review of machine learning in asset pricing, see the book of Nagel (2021).
One alternative to neural nets is the Kolmogorov–Arnold representation theorem (Kolmogorov 1956; Arnold 1957) introduced in finance by Borri et al. (2024). [^9] Borri et al. (2024), arXiv, April 11, 2024. Kolmogorov–Arnold remarkably shows that any nonlinear function can be written exactly as a combination of a finite sum of univariate functions composed with a link function. This formulation is thus a representation, not an approximation. Importantly, the early machine learning literature showed how to use in theory this representation to construct a shallower and smaller network (Hecht-Nielsen 1987, Girosi & Poggio 1989, Kurkova 1991). This network can be thought of as a much lower dimensional neural network where the activation functions of each neuron are being trained and adapt to the data. While theoretically this result is remarkable, [^10] In fact, Kolmogorov and Arnold construction solved Hilbert’s thirteenth problem, one of the major open problems of mathematics of the twentieth century. it is not clear whether this approach would work on actual data because of the complexity that the theory requires for the representation of any function, and in particular, on finance data, which is both relatively small and has a very low signal-to-noise ratio. Borri et al. (2024) is the first paper that shows the practical implementation and effectiveness of the Kolmogorov–Arnold architecture on actual data. Specifically, they show that this methodology enables construction of a single non-linear factor (KA factor) that captures a significant part of the variation in expected returns. [^11] Schmidt-Hieber (2021) is the first paper in the literature that studies quantitative properties of Kolmogorov–Arnold representation for multiple-layer neural networks and thus introduces the Kolmogorov–Arnold architecture. An early paper by Horowitz & Mammen (2007) discusses essentially a similar structure of general nonparametric regressions with an unknown link function. Subsequent to Borri et al. (2024), Liu et al. (2024) (arXiv, April 30, 2024) popularized Kolmogorov Arnold Networks (KAN) to the machine learning community and showed how significantly shallower neural nets built on Kolmogorov–Arnold architecture may, in some cases, outperform and bring interpretability in several science applications. Toscano et al. (2025) argue that KAN rely on computation expensive basis functions with a performance that degrades with high parameter counts, which might limit the use in these science applications. The main advantage of this approach is that the KA-factor can be thought of as a simple representation of a much more complicated and larger neural network and, in principle, captures all such networks, without having to consider a variety of neural networks specifications and the optimization algorithms.
Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion. However, the Taylor approximation to the higher order generates a very large set of factors, due to the need to account for all of the interactions of the higher-order terms in the Taylor expansion. Borri et al. (2025) develop the properties of a forward-selection Fama–MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns. The method proceeds sequentially, adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared. The main advantage of this approach is that it is a fully “glass box” model both in terms of determining the important factors and in the process to select them that allows to map features of the data to the economic factors and their higher order interactions. [^12] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian & Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017). See Bell et al. (2024) for an alternative “glass box” machine learning structure.
Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu et al. (2021). The test assets are the set of 50 portfolios formed using the smart beta strategies described in Fact 4 (see Tables 2 and 3), and the CARB strategy described in Fact 8 (see Table 6, Panel B). As with the classic CAPM, the C-1 model cannot account for the cross-section of average returns. By contrast, the C-4 model explains 47.3% of the cross-sectional variation, and satisfies the restriction of a zero intercept.
Table 5, Panel B, reports the result of a specification which augments the C-4 model with the single KA-factor based on the Kolmogorov-Arnold architecture of Borri et al. (2024). One can think of this factor as a stand-in that captures a variety of different neural nets. The cross-sectional R-squared increases by approximately 25 pp, and the intercept is not statistically different from zero at standard levels.
Table 5, Panel C, reports the results from the estimation of the FS-FMB (Borri et al. 2025). At each step, we augment the C-4 by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. Higher-order factors include the squared and cubed of the factors in the C-4 model, as well their pair-wise interactions up to order three. The procedure selects three higher-order factors. The first is CMOM cubed, resulting in an increase of the cross-sectional R-squared of approximately 15 pp. The importance of this factor is that it relates to the frequent and large jumps which are the distinct feature of cryptocurrency of Fact 5. The second and third selected factors are CMKT squared and the interaction between CMKT and CSIZE. Borri et al. (2025) show that the equity counterparts of these factors are important for the cross-section of equity returns. The cross-sectional R-squared of the FS-FMB procedure is 70%, a value very similar to that of the model including the KA-factor. In all models, the estimate of the intercept is not statistically different from zero.
These results show that the FS–FMB procedure recovers essentially the same nonlinear structure as the KA single factor. Moreover, since KA can represent any neural net, FS–FMB provides a simple, transparent, “open box” alternative to black-box models that is directly connected to the economic structure of the model and to observables characteristics. In short, rather than estimating a neural network (or studying a particular variant of the multitude of neural networks), one can estimate the FS–FMB specification and capture the relevant nonlinearities in a parsimonious and interpretable way without relying on millions of parameters.
Table 5: Cross-Sectional Performance
Panel A: Baseline Models
# Model Adj. R-squared $\alpha$ t-stat ($\alpha$)
1 C-1 0.046 0.012 1.437
2 C-4 0.473 0.005 0.766
Panel B: KA model
# Model Adj. R-squared $\alpha$ t-stat ($\alpha$)
1 C-4 & KA 0.723 0.011 1.494
Panel C: FS-FMB procedure
Step $h_{j}$ Adj. R-squared $\alpha$ t-stat ($\alpha$)
1 CMOM3 0.623 0.002 0.235
2 CMKT2 0.663 -0.001 -0.140
3 CMKT $\times$ CSIZE 0.702 0.001 0.138
This table reports the results from the second step of the FMB method. Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu et al. (2022). The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic. The t-statistics is based on Newey-West corrected standard errors. Panel B refers to the C-4 model augmented by the single KA-factor, estimated with the KA procedure of Borri et al. (2024). Panel C refers to the FS-FMB procedure. At each step, we augment the C-4 model by one additional higher-order factor ($h_{j}$) which maximizes the adjusted cross-sectional R-squared. The method stops when the improvement in the R-squared in a given step is smaller than 1 pp.
3.7 Fact 7: In crypto, the (block)chain drives the gain
Fact 6: Blockchain economics and information significantly affects cryptocurrency prices, on-chain user adoption accounts for about 8% of the variation in cryptocurrency returns.
Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption. Classic work has analyzed how networks form and evolve (Jackson & Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet & Tirole, 2003), and how shocks propagate through financial networks (Allen & Gale, 2000). A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010, Acemoglu et al. 2011).
One of the defining features of blockchain and one of the important innovations that it delivers is that blockchains create a continuously updated, public, and verifiable record of economic activity. Blockchains provide an observable, real-time map of user adoption and economic interactions. This makes them particularly useful for studying how information absorption affects asset values. As Metcalfe’s law put it, “the value of a network grows as the square of the number of its users.”
Empirically, Liu et al. (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices. Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns. In time-series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
Liu et al. (2021) also examine how markets absorb this information and what conditions amplify its relevance. Event-time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre-announcement behavior and no post-announcement drift—in contrast to equities. The result gives evidence that investors fully incorporate the information at the weekly frequency.
Lastly, Liu et al. (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in Fact 4.
3.8 Fact 8: Young cryptocurrency markets, old inefficiencies.
Fact 8: The young cryptocurrency markets are still inefficient. But the origins are canonical: market segmentation and restrictions on capital flows. Cross-exchange cryptocurrency “arbitrage” is no free lunch.
A central concept in finance is the law of one price—identical securities should sell for identical prices. The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978, Harrison & Kreps 1979). Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont & Thaler 2003), limits to arbitrage and market segmentation (Shleifer & Vishny 1997), or previously overlooked risks (Cochrane 2002).
One of the main features of cryptocurrency is that it is possible to trade coins simultaneously on multiple exchanges and against a number of fiat and crypto pairs. For example, an investor could buy Bitcoin on Binance or Kraken, two prominent centralized exchanges, against fiat currencies such as the U.S. dollar or the euro, or even against another cryptocurrency such as Ethereum or Tether, a stablecoin. Makarov & Schoar (2020) and Borri & Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs. We revisit this evidence on updated data for 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Bitcoin price discounts are computed each week relative to the dollar price on Kraken (the benchmark exchange). All quotes are converted to U.S. dollars using daily spot FX rates. In the 2014–2025 sample, the 10th and 90th percentiles of Bitcoin price discounts are –1.8% and 3.7%, respectively. These figures remain virtually unchanged in the post-2020 period (–2.1% and 3.5%). Moreover, discounts are highly persistent, with a half-life of roughly 1.1 weeks in both samples. These deviations—classic examples of market inefficiency and violations of the law of one price—are results of market segmentation induced by capital controls and idiosyncratic domestic shocks that prevent arbitrage. [^13] Barbon & Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
Table 6, Panel A, reports the properties of an “arbitrage” (CARB) strategy that attempts to exploit these price discrepancies. Each week, we sort BTC–fiat pairs by their discount and form five portfolios. The CARB return for a given pair is defined by buying Bitcoin on Kraken in week $t$ and selling it on the target pair in week t+1t{+}1; portfolio returns are the simple average across constituent pairs. Using weekly formation and holding periods substantially alleviates execution and timing concerns. In practice, capital controls, withdrawal limits, transaction costs, settlement latency, and venue-specific frictions prevent the cross-market trades needed to enforce the law of one price. Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly. Specifically, Borri & Shakhnov (2022) shows that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe. That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer & Vishny (1997).
As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades. [^14] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman Fried. However, Borri & Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish łoty, Swiss franc, British pound, and U.S. dollar). These are the pairs that are realistically investable. Table 6, Panel B, documents that excess returns remain statistically significant but are orders of magnitude smaller: the long–short weekly return declines from about $0.68$ to roughly $0.038$ in the full sample, and measures around $0.041$ in the post-2020 sample. In other words, where the strategy is actually implementable, the apparent “free lunch” largely vanishes. These returns are not pure arbitrage—they are compensation for risk. Borri & Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu et al. (2023), is the most visible manifestation of the results described in this fact.
Table 6: Cryptocurrency “Arbitrage” Strategy
Panel A: All exchanges
Quintiles
1 2 3 4 5 5–1
Full sample
CARB Low High
Mean -0.011** 0.009* 0.012** 0.017*** 0.674*** 0.685***
tt(Mean) (-2.22) (1.73) (2.43) (3.38) (4.39) (4.46)
Post-2020 sample
CARB Low High
Mean -0.014** 0.006 0.010* 0.015*** 0.940*** 0.954***
tt(Mean) (-2.43) (1.01) (1.76) (2.75) (3.90) (3.97)
Panel B: Exchanges in open markets
Quintiles
1 2 3 4 5 5–1
Full sample
CARB Low High
Mean -0.002 0.007 0.010* 0.014*** 0.036*** 0.038***
tt(Mean) (-0.39) (1.36) (1.94) (2.71) (7.66) (12.45)
Post-2020 sample
CARB Low High
Mean -0.008 0.005 0.008 0.012** 0.033*** 0.041***
tt(Mean) (-1.36) (0.75) (1.26) (2.13) (6.06) (9.90)
Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for the “arbitrage” (CARB) investment strategy that sorts Bitcoin trading pairs by their price deviation from a benchmark Bitcoin price on Kraken. Each week, pairs are sorted into five portfolios. The long-short portfolio is long portfolio 5 and short portfolio 1. The mean returns are the time-series averages of weekly equally-weighted portfolio excess returns. Panel A uses the full sample of 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies. Panel B restricts to 226 BTC–fiat pairs on 96 exchanges against 11 fiat currencies, which are the most liquid fiat currencies on the most reputable exchanges in jurisdictions with minimal capital controls as in Borri & Shakhnov (2022). The post-2020 sample is the sample since Jan 1, 2020. $t$-statistics in parentheses are Newey-West adjusted. *, **, and *** denote significance at the 10%, 5% and 1% levels.
3.9 Fact 9: When the funding dries up, we finally learn the worth of futures
Fact 9: Cryptocurrency carry trade, the defining characteristics of the futures market, used to be highly profitable, but profitability has compressed sharply since 2024.
Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973, Merton 1973, Black 1976).
There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs. Among all derivatives, perpetual futures are the most common derivative contract in cryptocurrency markets. For example, Schmeling et al. (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample. Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot. Ackerer et al. (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures. The crypto-carry trade strategy of Schmeling et al. (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
Figure 2 plots the evolution over time of one dollar invested in the crypto-carry strategy (black line, left y-axis) and in the spot Bitcoin market (red line, right y-axis). Over the full sample, which goes from 2020 to 2025, the annualized Sharpe ratio of the cryptocurrency carry is 6.45. Beginning in 2024, the Sharpe ratio falls to 4.06, and it turns negative in 2025. The profit from the cryptocurrency carry strategy is mostly driven by the funding rate, which in the full sample has a mean return of approximately 8% with a low volatility of 0.8%.
This shift in the profitability of the crypto-carry trade strategy has important implications for the crypto-industry. A growing set of crypto “yield” products—most prominently delta-neutral, funding-rate-harvesting structures (e.g., synthetic dollar projects that combine spot holdings with short perpetuals)—derive an important part of their returns from the funding rate. A prominent example is Ethena.fi, whose stablecoin uses a delta-neutral construction and relies on perpetual-funding payments. As of September 2025, Ethena’s stablecoin reached a market capitalization of about \$14 billion, with its governance token ENA valued at around \$4 billion. If funding premia compress or become more volatile, the economics of these strategies deteriorate and the expected yields decline. The lower profitability of the crypto-carry trade shows that funding-rate premia are neither guaranteed nor permanent and, even more importantly, raises difficult questions about the long-run sustainability of yield products built on them.
Figure 2: Carry Trade Cumulative Returns
The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis). The Carry Trade is defined as in Schmeling et al. (2023). The sample goes from August 1, 2020 to May 31, 2025.
3.10 Fact 10: Growing up with supervision—regulation and oversight strengthen cryptocurrency markets
Fact 10: As the market matures, there is growing disclosure and transparency in cryptocurrencies.
A fundamental insight of information economics is that markets with asymmetric information can fail to allocate resources efficiently. When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970). Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983, Dye 1985, Healy & Palepu 2001, Verrecchia 2001). However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001, Ma & Shleifer 2025). Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information. Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts & Zimmerman 1983).
Information asymmetry remains pervasive in cryptocurrency markets, though regulation and disclosure have increased in recent years. On the corporate side, public firms with material cryptocurrency exposure are now required to disclose their cryptocurrency holdings and risks. Importantly, the introduction of Accounting Standards Update (ASU) 2023-08 by the Financial Accounting Standards Board provides a coherent framework for the accounting of crypto-assets, which formally requires fair-value accounting treatment (FASB 2023). Recent accounting research emphasizes measurement reliability and disclosure consistency for digital assets (Anderson et al. 2025).
On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021, Liu et al. 2025, Du et al. 2025). Howell et al. (2020) and Bourveau et al. (2022) show the importance of voluntary disclosures in the initial coin offering stage. Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al. 2022, Bourveau et al. 2024).
In July 2025, Congress enacted the Guiding and Establishing National Innovation for U.S. Stablecoins Act (GENIUS Act), the first comprehensive federal framework for digital asset oversight in the United States. The Act focuses primarily on payment stablecoins, aiming to improve their transparency and stability. It mandates monthly public disclosures of reserve compositions of stablecoins, and explicitly prohibits stablecoin companies from making misleading claims that their tokens are backed by the U.S. government, federally insured, or legal tender.
The recent regulations and market developments show that greater transparency in the cryptocurrency space is not only increasingly required by the regulators but is also valued by the market. The cryptocurrency market is starting to increasingly implement and value more rigorous accounting principles and slowly converging to the accounting practices of traditional asset classes, for which reliable, comparable, and verifiable disclosure is a major and perhaps the most important determinant of investor confidence and growth of an asset class.
4 ADDITIONAL DISCUSSIONS
The ten facts summarize key developments in cryptocurrencies through the lens of empirical asset pricing. In this section, we provide additional discussions on emerging topics, including retail investor trading in the cryptocurrency market, the non-fungible token market, and studies on the dark side of cryptocurrency usage.
Recent research increasingly examines individual investor beliefs and behavior in cryptocurrency markets and compares and contrast them with those in traditional investments. Kogan et al. (2024) show that the trading behaviors of retail investors are drastically different between cryptocurrencies and traditional assets, such as stocks and gold. In particular, retail investors are contrarian in stocks and gold, but trend-chasing in cryptocurrencies. Kogan et al. (2024) conjecture that the trend-chasing behavior is due to cryptocurrency price changes affecting the likelihood of future widespread adoption, which leads retail investors to update their price expectations. Chan et al. (2020) use data from a medium-sized cryptocurrency exchange and document that investor portfolios are typically small in value, highly concentrated in a few coins, and characterized by short holding periods. Weber et al. (2023) complement this evidence using repeated large-scale surveys of U.S. households, showing that cryptocurrency investors differ systematically from other investors. Alvarez et al. (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption. Han et al. (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets. Finally, Benetton & Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies. They find that individual beliefs of cryptocurrency are different and are critical in explaining cryptocurrency valuation.
The bulk of this review is about fungible tokens. A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs). Goetzmann & Nozari (2022) is the first to construct repeated sales return indices of NFTs and examine their properties. Borri et al. (2022a) provide a comprehensive analysis of the NFT market. They highlight the importance of visual features in cryptocurrency valuation and how it leads to a conspicuous consumption effect in the market. Furthermore, Borri et al. (2022a) shows that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market. Huang & Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing. Oh et al. (2025) document preference-driven herding in NFT primary market sales.
A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years. Foley et al. (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley et al. (2019) are largely driven by double-counting. Griffin & Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation. Amiram et al. (2022) show that abnormal spikes in on-chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, and Cong et al. (2025) provide an anatomy of crypto-enabled crimes. Li et al. (2025) document pervasive cryptocurrency pump-and-dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre-event run-ups consistent with insider gains at outsiders’ expense.
5 CONCLUSIONS
In the 16 years since Nakamoto, cryptocurrencies have reached a certain stage of maturity: they can now be studied using the same measurement, data quality, and research rigor applied to other asset classes. The market is now worth more than four trillion dollars. Major financial firms are using cryptocurrency in their businesses.
We show that there are important similarities between cryptocurrency markets and traditional asset markets—Sharpe ratios are broadly comparable, and a small number of factors and their higher orders summarize the cross-section of returns. On the other hand, cryptocurrency also has its own character—frequent jumps and blockchain information driving prices. The cryptocurrency market has become more mature and is emerging as an asset class with its own distinct characteristics that still, however, needs more data, research, supervision, enforcement, and a better legal and regulatory framework.
Given the maturation of cryptocurrency into an asset class, this review shows that cryptocurrency should be studied the same way we study any asset. The right way to understand it is exactly through traditional finance tools—empirical asset pricing. The ten facts we present summarize the key takeaways from this view.
DISCLOSURE STATEMENT
The authors are not aware of any affiliations, memberships, funding, or financial holdings that might be perceived as affecting the objectivity of this review. This review is for academic purposes only and should not be construed as investment or legal advice.

### v3

@@
 Abstract
-Cryptocurrencies are coming of age.
-We organize empirical regularities into ten stylized facts and analyze cryptocurrency through the lens of empirical asset pricing.
-We find important similarities with traditional markets—risk-adjusted performance is broadly comparable, and the cross-section of returns can be summarized by a small set of factors.
+We organize existing empirical regularities of cryptocurrencies into seven stylized facts and analyze cryptocurrencies through the lens of empirical asset pricing.
+We find important similarities with traditional markets—risk-adjusted performance so far is broadly comparable, and the cross-section of returns can be summarized by a small set of factors.
 However, cryptocurrency also has its own distinct character: jumps are frequent and large, and blockchain information helps drive prices.
-This common set of facts provides evidence that cryptocurrency is emerging as an investable asset class.
+This common set of stylized facts provides evidence that cryptocurrency is emerging as an investable asset class.
+Additionally, we discuss potential data quality issues and possible changes in future regulations and the cryptocurrency environment.
 1 INTRODUCTION
-This survey establishes ten key empirical regularities of cryptocurrencies motivated by the view that cryptocurrency is an asset and thus can and should be analyzed from the asset-pricing point of view and, more specifically, from the cryptocurrency asset pricing view.
+This survey establishes a set of seven key empirical regularities of cryptocurrencies motivated by the view that cryptocurrency is an asset and thus can and should be analyzed from the asset-pricing point of view and, more specifically, from the cryptocurrency asset pricing view.
 Although the literature is relatively young—cryptocurrencies trace back to Nakamoto (2008)’s short paper on Bitcoin—it is now at the point of maturity where it is useful to establish a core of stylized facts in cryptocurrency markets that can be used for both empirical and theoretical work.
@@
 When the paper of Nakamoto (2008) introduced a permissionless blockchain that provides public record-keeping without a central authority, financed and governed by a native token, the canonical literature on money had reached a degree of maturity.
-Permissionless ledgers can be viewed as technologies that supply public memory, with token systems financing and governing that record—keeping, and providing histories, communication, and enforcement mechanisms that are required for credit.
+Permissionless ledgers can be viewed as technologies that supply public memory, with token systems financing and governing that record-keeping, and providing histories, communication, and enforcement mechanisms that are required for credit.
 With the rise of cryptocurrencies, these new ideas began to emerge primarily from practice more than theory and reopened classic monetary questions—record-keeping, commitment, enforcement—but now viewed from the point of view of blockchain economics.
@@
 The paper of Liu & Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing.
-A sizable literature uses an asset pricing lens to study risk exposures, cross-sectional return patterns, cash-flow-like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020, Liu et al.
-2022, Alvarez et al.
-2023, Sockin & Xiong 2023a, von Luckner et al.
-2023, Kogan et al.
+A sizable literature uses an asset pricing lens to study risk exposures, cross‑sectional return patterns, cash‑flow‑like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020; Liu, Tsyvinski & Wu 2022; Alvarez, Argente & Van Patten 2023; Sockin & Xiong 2023a; von Luckner, Reinhart & Rogoff 2023; Kogan et al.
 2024).
@@
 When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system.
-[^1] Foley et al.
-(2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations.
-It is worth noting that several important financial innovations, such as the limited-liability company and mortgage-backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
-Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018, Leshno & Strack 2020, Biais et al.
-2019, Kogan et al.
-2022, Brunnermeier et al.
-2019, Benigno et al.
-2022, Capponi et al.
-2025, Jermann & Xiang 2025), regulation (Cong & He 2019,Townsend 2020, Auer et al.
-2025), digital asset pricing and valuation (Athey et al.
-2016, Cong et al.
-2021b, Biais et al.
-2023), transaction costs (Easley et al.
-2019, Cong et al.
-2021a, Basu et al.
+[^1] See, Foley, Karlsen & Putniņš (2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations.
+It is worth noting that several important financial innovations, such as the limited‑liability company and mortgage‑backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
+Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018; Leshno & Strack 2020; Biais et al.
+2019; Kogan, Fanti & Viswanath 2022; Brunnermeier, James & Landau 2019; Benigno, Schilling & Uhlig 2022; Capponi, Jia & Wang 2025; Jermann & Xiang 2025), regulation (Cong & He 2019; Townsend 2020; Auer, Monnet & Shin 2025), digital asset pricing and valuation (Athey et al.
+2016; Cong, Li & Wang 2021b; Biais et al.
+2023a), transaction costs (Easley, O’Hara & Basu 2019; Cong, He & Li 2021a; Basu et al.
 2023), central bank money (Auer et al.
-2022, Whited et al.
-2023, Schilling et al.
-2024), securitization (Cong et al.
-2022, Sockin & Xiong 2023a, Rogoff & You 2023, Goldstein et al.
-2024), safe assets (Ma et al.
-2025), international capital flow (von Luckner et al.
-2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al.
+2022; Whited, Wu & Xiao 2023; Schilling, Fernández-Villaverde & Uhlig 2024), securitization (Cong, Li & Wang 2022; Sockin & Xiong 2023a; Rogoff & You 2023; Goldstein, Gupta & Sverchkov 2024), safe assets (Ma, Zeng & Zhang 2025), international capital flow (von Luckner, Reinhart & Rogoff 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al.
 2024).
-We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of ten stylized empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes.
-These empirical facts are also useful to evaluate and revisit both the classic and the new theories of money.
-Ten Facts in Cryptocurrency Asset Pricing and Empirical Finance.
-Facts 1 and 2 overview the core asset-pricing characteristics of cryptocurrency markets—the risk-return trade-off and the diversification benefits.
-Facts 3 and 4 discuss the structure of returns in cryptocurrency—factors, or “smart-beta”, which explain the cross-section of returns and optimal portfolio shares.
-Fact 5 shows the importance of large, discrete price movements—jumps—and the importance of extreme events for understanding cryptocurrency risk and valuation.
-Fact 6 shows that the higher-order terms of a parsimonious four-factor model account for a large fraction of the cross-section of returns and provide a “glass box” alternative to the “black box” neural net models with interpretable economic factor structure.
-Fact 7 shows the importance of the blockchain structure and network economy phenomena for cryptocurrency prices—on-chain user adoption influences cryptocurrency returns.
-Facts 8 documents large and persistent market inefficiencies due to frictions that limit arbitrage, and discuss risks for arbitrageurs.
-Fact 9 is about the cryptocurrency futures market and funding premia, and their evolution.
-Finally, Fact 10 is about the importance of regulation, disclosure, and transparency.
-In sum, these ten facts can be viewed as providing evidence that cryptocurrency, and more broadly, digital assets are coming of age as an emerging investable asset class.
+We synthesize the rapidly growing empirical finance of cryptocurrencies into a set of seven stylized facts—empirical regularities that show how cryptocurrencies behave as assets, how their markets function, and what distinguishes them from traditional asset classes.
+These stylized facts help evaluate and revisit both classic and new theories of money, while remaining conditional on the available data and institutional environment to date and therefore potentially evolving as market structure, regulation, and adoption change.
+[^2] For complementary surveys on blockchain architecture, decentralized finance and central bank digital currencies (CBDC), see Kose et al.
+(2022); Makarov et al.
+(2022); Kose et al.
+(2023); Biais et al.
+(2023b); Adrian & Mancini-Griffoli (2021); Infante et al.
+(2023).
+Seven Stylized Facts in Cryptocurrency Asset Pricing and Empirical Finance.
+Stylized fact 1 overviews the core asset-pricing characteristics of cryptocurrency markets—the risk-return trade-off and the diversification benefits.
+Stylized fact 2 discusses the structure of returns in cryptocurrency—factors, or “smart-beta”, which explain the cross-section of returns.
+Stylized fact 3 shows that the higher-order terms of a parsimonious four-factor model account for a large fraction of the cross-section of returns and provide a “glass box” alternative to the “black box” neural net models with interpretable economic factor structure.
+Stylized fact 4 shows the importance of the blockchain structure and network economy phenomena for cryptocurrency prices—on-chain user adoption influences cryptocurrency returns.
+Stylized fact 5 documents large and persistent market inefficiencies due to frictions that limit arbitrage, and discusses risks for arbitrageurs.
+Stylized fact 6 is about the cryptocurrency futures market and funding premia, and their evolution.
+Finally, stylized fact 7 is about the importance of regulation, disclosure, and transparency.
+In sum, these seven facts can be viewed as providing evidence that cryptocurrency, and more broadly, digital assets are coming of age as an emerging investable asset class.
 2 DATA
@@
 The final dataset comprises 16,468 unique coins and approximately 9.5 million records.
-Second, for the construction of the crypto carry strategy of Schmeling et al.
-(2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025.
-The data on new addresses are from Liu et al.
-(2021).
+Second, for the construction of the crypto carry strategy of Schmeling, Schrimpf & Todorov (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025.
+The data on new addresses are from Liu, Tsyvinski & Wu (2021).
 Moreover, for the construction of the crypto arbitrage strategy of Borri & Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025.
-Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.222Finally, we obtain price indices for the U.S.
-stock market, U.S.
-corporate and government bonds, a broad commodity index, gold, the U.S.
+We convert these quotes into U.S.
+dollars using spot FX rates from Bloomberg.
+Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.333Finally, we obtain price indices for the MSCI global stock market, global high-yield and government bonds, a broad commodity index, gold, the U.S.
 dollar index, the Dow Jones US Real Estate Index, the U.S.
 constant maturity 10-year nominal and real bond yields from Bloomberg.
-The corresponding Bloomberg tickers are SPX Index, LUACTRUU Index, LUATTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
+The corresponding Bloomberg tickers are MXWO Index, LG30TRUU Index, LGTRTRUU Index, BCOM Index, XAU Curncy, DXY Curncy, DJUSRE Index, USGG10YR Index and GTII10 Govt.
 Unless otherwise noted, we aggregate daily closing prices to construct weekly coin returns.
+[^4] CoinGecko daily prices are timestamped at 00:00 UTC (19:00 ET), whereas traditional asset prices from Bloomberg reflect end-of-day timestamps that vary by market and instrument (typically between 16:00 and 19:00 ET).
+Since our baseline analysis uses weekly close-to-close returns, any intraday timing mismatch is negligible.
+Crypto prices are quoted directly in USD, so no additional currency conversion is required.
+This review follows the dominant practice of using USD-denominated prices, which reflects the phenomenon that cryptocurrency trading is largely dominated by USD-linked quote currencies.
+We note that alternative numeraire or crypto-to-crypto return measures may reveal additional dimensions that are not captured by USD-based trades, which is a potentially important direction for future research.
 Each year is divided into 52 weeks: the first week contains the year’s first seven days, the next 50 weeks consist of seven days each, and the final week contains the remaining days of the year.
@@
 Louis.
-Figure 1 summarizes the time-series evolution of cryptocurrency prices and market capitalization.
+An important caveat concerns data integrity.
+A strand of the empirical cryptocurrency literature argues that parts of the historical crypto data—especially in early years—may reflect manipulation or unreliable reporting, including alleged price manipulation on Mt.
+Gox in 2013 (Gandal et al., 2018), inflated or “artificial” trading volume reported by some exchanges (Makarov & Schoar, 2020), and exchange-managed wash trading concentrated on “unregulated” exchanges and largely absent on “regulated” exchanges such as Bitstamp, Coinbase, and Gemini (Cong et al., 2023b), as well as trade reordering/front-running on decentralized exchanges (Daian et al., 2019).
+At the same time, on-chain evidence points to substantial cross-exchange flows: analyzing the Bitcoin blockchain, Makarov & Schoar (2021) show that a sizable share of volume to exchanges originates from other exchanges or large market participants, consistent with arbitrage activity that tends to limit persistent price deviations across major venues.
+While these concerns cannot be fully eliminated, we mitigate them in several ways.
+First, our CoinGecko sample starts after the alleged Mt.
+Gox manipulation episode in 2013.
+Second, our pre-/post-2020 split places emphasis on the most recent period, and the post-2020 subsample begins after the main evidence documented in Makarov & Schoar (2020) and Cong et al.
+(2023b).
+Third, the CoinGecko prices we use are volume-weighted averages across a curated set of large spot exchanges and, per CoinGecko’s documentation, [^5] See, https://www.coingecko.com/en/methodology.
+are constructed using algorithms designed to detect and ignore outliers; as of the end of our sample these include the following exchanges: Binance, Coinbase, Kraken, Crypto.com, Bitfinex, Bitstamp, Gemini, and Huobi.
+Finally, our market-capitalization screens and data filters are designed to downweight or exclude small, thinly traded coins where data quality issues are more likely to be severe.
+Figure 1 summarizes the time-series evolution of cryptocurrency prices and market capitalization through September 2025.
 A one-dollar investment at the start of the sample grows to about \$160 in Bitcoin and \$80 in the overall market.
 There is substantial variation over time.
-Market capitalization reaches nearly \$4 trillion, with Bitcoin and Ethereum accounting for approximately 58% and 12% of the total at the end of the sample.
+Market capitalization reaches about \$4 trillion, with Bitcoin and Ethereum accounting for approximately 58% and 12% of the total at the end of the sample.
+[^6] Since the end of our sample, aggregate cryptocurrency market capitalization has declined; as of late January 2026 it is about \$2.8 trillion.
 (a) Return
@@
 Data end on September 6, 2025.
-3 THE TEN FACTS OF CRYPTOCURRENCY
-We organize the main discussions of this review into the ten key stylized empirical facts on cryptocurrency markets.
-3.1 Fact 1: High return, high volatility—normal Sharpe ratio
-Fact 1: The cryptocurrency market has an order of magnitude higher returns and volatility than equities, but a broadly similar Sharpe ratio.
+3 THE SEVEN STYLIZED FACTS OF CRYPTOCURRENCY
+We organize the main discussions of this review into seven key empirical regularities, or “stylized facts,” on cryptocurrency markets.
+3.1 Stylized fact 1: High return, high volatility—normal Sharpe ratio so far, but rising comovement with equity
+Stylized fact 1: The cryptocurrency market has an order of magnitude higher returns and volatility than equities, but a broadly similar Sharpe ratio; since 2020, it has also become substantially more correlated with equities.
 Popular discussion of cryptocurrencies tends to focus either on their dramatic average returns or on dramatic volatility.
@@
 Liu & Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable.
-Fact 1 revisits and confirms this finding in a longer and larger dataset up to September 2025.
+Stylized fact 1 revisits and confirms this finding in a longer dataset through September 2025.777These performance measures are computed from price indices and thus reflect the return to the underlying asset, not the return to a particular investor after custody risk.
+Episodes such as exchange cybersecurity breaches or bankruptcies (e.g., Mt.
+Gox and FTX) can generate additional losses that are not reflected in spot price series.
 Table 1, Panel A, reports summary statistics for the sample beginning in 2014 and ending in September 2025, as well as for the samples pre-2020 and post-2020.
-We compare returns on the aggregate coin market, a large-cap index of the top 100 cryptocurrencies (Coin 100), Bitcoin, and the aggregate U.S.
-stock market.
-The average returns and volatilities of all cryptocurrency return series are an order of magnitude greater than those of the stock market across all sample periods.
-For instance, average weekly returns exceed 1% for all cryptocurrency indices, compared with less than 0.3% for the aggregate stock market.
-However, the weekly Sharpe ratios are similar across cryptocurrencies and equities.
-This, confirms that the Sharpe ratio calculated by Liu & Tsyvinski (2021) holds in the most recent sample.
-Table 1, Panel A, reports maximum drawdowns and run-ups for the four return indices.
-Drawdowns are important as they provide a measure of tail-risk not fully reflected by standard deviation.
-Drawdowns and run-ups are substantially larger for cryptocurrencies than for equities.
-Importantly, for all three cryptocurrency indices, maximum drawdowns and run-ups are smaller in magnitude in the post-2020 sample than in the pre-2020 sample, while the reverse holds for equities.
-This pattern can also be observed in the standard deviations of returns—they decline for cryptocurrencies and rise for equities in the more recent period.
+We compare returns on the aggregate coin market, a large-cap index of the top 100 cryptocurrencies (Coin 100), Bitcoin, and the aggregate global equity market.
+Across all three cryptocurrency indices, average weekly returns exceed 1% and volatilities are about an order of magnitude larger than for global equities; however, weekly Sharpe ratios are similar across cryptocurrencies and equities.
+Panel A also reports drawdowns and run-ups, which are substantially larger for cryptocurrencies than for equities.
+Importantly, drawdowns and return volatility for cryptocurrencies are smaller in the post-2020 sample than pre-2020, while the reverse holds for equities.
+A complementary change concerns diversification.
+Table 1, Panel B, reports correlations between the coin market return and major asset classes.
+In the pre-2020 period, correlations with traditional assets are small and statistically insignificant, consistent with Liu & Tsyvinski (2021).
+In the post-2020 period, correlations rise sharply—most notably with global equities, increasing from 1% pre-2020 to 39% post-2020—and the coin market also becomes significantly correlated with commodities (27%), high-yield bonds (24%), government bonds (13%), and gold (18%).
+We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S.
+Treasury bonds with 10-year maturity.
+Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years.
+The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong, Fang & Lin (2023) and Jiang et al.
+(2025).
+Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization.
+The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market.
+[^8] Results are similar using a broad U.S.
+stock market index such as the S&P 500 (2% pre-2020; 37% post-2020), while correlations with U.S.
+government bonds (rather than global government bonds) remain insignificant in the post-2020 period.
+These correlation patterns matter directly for portfolio choice.
+In mean–variance settings, an asset’s contribution to diversification depends on its covariance with the rest of the portfolio rather than its standalone volatility (Markowitz, 1952).
+Consistent with the portfolio results in Liu & Tsyvinski (2021) and related Bayesian approaches (Duchin et al., 2025), a simple Black–Litterman calibration using post-2020 moments still implies non-trivial but modest allocations to the aggregate crypto market—about 3.1% for a representative retail portfolio and 5.5% for an institutional portfolio with a broader opportunity set.
+We emphasize that these magnitudes are back-of-the-envelope and depend on strong assumptions about expected returns, correlations, and investability (liquidity, custody, and regulatory constraints), and should therefore be interpreted as illustrative upper bounds on diversification benefits rather than investment advice.
 Table 1: Summary statistics
@@
 Bitcoin 1.33 10.12 3.24 0.13 0.33 2.16 -83.62 208.94
-Stock 0.27 2.18 3.01 0.11 -0.66 3.95 -27.22 17.55
+Stock 0.17 2.10 2.06 0.07 -0.74 4.79 -28.04 16.00
 Pre-2020 sample
@@
 Bitcoin 1.35 11.28 2.12 0.12 0.31 1.46 -83.62 208.94
-Stock 0.24 1.79 2.35 0.12 -0.68 2.54 -17.09 10.67
+Stock 0.13 1.74 1.29 0.06 -0.59 1.69 -18.57 10.61
 Post-2020 sample
@@
 Bitcoin 1.30 8.75 2.56 0.14 0.32 3.03 -74.94 95.10
-Stock 0.29 2.53 2.00 0.10 -0.63 3.47 -27.22 17.55
+Stock 0.23 2.42 1.60 0.07 -0.80 4.76 -28.04 16.00
 Panel B. Coin market return correlations
@@
 (1) (2) (3) (4) (5) (6) (7)
-Full 0.20 0.08 0.03 0.14 0.12 -0.06 0.17
+Full 0.20 0.14 0.09 0.14 0.12 -0.06 0.17
 (4.93) (3.51) (2.10) (3.60) (2.87) (-1.49) (4.35)
-Pre 2020 0.02 0.04 0.02 0.03 0.06 -0.03 0.03
+Pre 2020 0.01 0.03 0.05 0.03 0.06 -0.03 0.03
 (0.15) (0.48) (0.88) (0.51) (1.13) (-0.48) (0.46)
-Post 2020 0.37 0.13 0.04 0.27 0.18 -0.10 0.31
+Post 2019 0.39 0.24 0.13 0.27 0.18 -0.10 0.31
 (7.31) (4.28) (2.20) (4.74) (3.10) (-1.80) (5.66)
 Notes: This table summarizes the main properties of coin market return.
-Panel A reports weekly summary statistics of the coin market index and compares them with the value-weighted index of the top 100 coins by market capitalization (Coin 100), Bitcoin and U.S.
-stock market returns.
+Panel A reports weekly summary statistics of the coin market index and compares them with the value-weighted index of the top 100 coins by market capitalization (Coin 100), Bitcoin and global stock market returns.
 Reported measures include the mean, standard deviation, $t$-statistic, Sharpe ratio, skewness, kurtosis, maximum drawdown and run-up.
-Panel B presents pairwise correlation coefficients between coin market returns and returns on the U.S.
-stock market, corporate bonds, government bonds, a commodity index, gold, the dollar index, and the changes in inflation expectations.
+Panel B presents pairwise correlation coefficients between coin market returns and returns on the global stock market, global high-yield bonds, global government bonds, a commodity index, gold, the dollar index, and the changes in inflation expectations.
 $t$-statistics are reported in parentheses.
@@
 The pre-2020 sample is the sample before Jan 1, 2020, and the post-2020 sample is the sample since Jan 1, 2020.
-3.2 Fact 2: Cryptocurrency is a distinct asset: it moves largely on its own—but increasingly with equities, gold and commodities
-Fact 2: The aggregate cryptocurrency market remains largely uncorrelated with most traditional asset classes, but its correlation with U.S.
-equities has increased sharply—from 2% before 2020 to 37% after 2020.
-How cryptocurrencies co-move with traditional asset classes is a first-order question for asset pricing and portfolio choice (e.g., Markowitz 1952).
-Correlations are an important determinant of diversification benefits and show whether cryptocurrency risk is distinct to that market or simply represent familiar risks from other markets.
-We revisit the cryptocurrency aggregate correlation structure using updated data, contrasting the pre-2020 and post-2020 return correlation.
-One of the main results of Liu & Tsyvinski (2021) is that aggregate cryptocurrency returns have insignificant exposures to the returns of traditional asset classes, such as equities, bonds, commodities, and foreign exchanges.
-In other words, the correlation between the aggregate cryptocurrency returns and those of traditional asset classes are low.
-Fact 2 revisits this finding.
-We show that there is a rising correlation with the U.S.
-stock market.
-Table 1, Panel B, reports correlations of aggregate cryptocurrency returns with those of major asset classes.
-In the pre-2020 sample, correlations with equities, corporate bonds, government bonds, commodities, gold, and the dollar are all statistically insignificant, confirming the earlier findings of Liu & Tsyvinski (2021).
-In contrast, in the post-2020 sample, aggregate cryptocurrency returns become significantly correlated with equities (37%), commodities (27%), gold (18%), and corporate bonds (13%) at the 5% level.
-The correlation with the dollar index is negative (-10%) and significant at the 10% level.
-Government bonds remain the only asset class with no significant link.
-The most striking change is the correlation with the U.S.
-stock market, which rises from only 2% before 2020 to 37% afterwards.
-This substantial increase is an important indicator that cryptocurrencies have become more integrated with equity markets.
-We also document a significant post-2020 correlation with changes in expected inflation (31%), measured as the first difference in the yield differential between nominal and inflation-indexed U.S.
-Treasury bonds with 10-year maturity.
-Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years.
-In our view, this link with expected inflation is important, as it provides empirical evidence on the original motivation of cryptocurrencies as assets designed to hedge against the inflationary risks of government printing money.
-The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong et al.
-(2023) and Jiang et al.
-(2025).
-Several mechanisms are proposed to explain this phenomenon, including monetary policy, retail investor trading, and financialization.
-The literature has yet to reach a consensus on the cause of this important shift in the return dynamics of the cryptocurrency market.
-3.3 Fact 3: Significant diversification benefits from adding small amount of cryptocurrencies in investment portfolios
-Fact 3: Cryptocurrency idiosyncratic risk can be significantly reduced by constructing optimal mean-variance asset portfolios that include cryptocurrencies—the inclusion of cryptocurrencies enlarges the mean-variance frontier.
-Optimal portfolio allocates 3.1% cryptocurrency for retail investors and 5.5% for institutional investors.
-Modern finance theory, starting from Markowitz (1952), emphasizes the central insight of diversification: an asset’s covariance with the investor’s portfolio matters more than its volatility.
-Low correlations push the efficient frontier outward, and volatile assets can add value if their expected returns compensate for marginal risk.
-Since Fact 1 establishes that cryptocurrencies have low correlations with stocks and bonds, they can be effective risk diversifiers under standard portfolio-theory frameworks.
-Liu & Tsyvinski (2021) show that cryptocurrencies can improve portfolio efficiency because of their historically low correlations with traditional asset classes and comparatively high expected returns.
-Using the Black–Litterman portfolio allocation model (Black & Litterman 1990) to incorporate investor views, Liu & Tsyvinski (2021) estimate that a representative investor would optimally allocate between 1% and 6% of wealth to cryptocurrencies, depending on the strength of those views.
-A complementary view comes from Bayesian portfolio theory.
-Duchin et al.
-(2025) obtain similar optimal weights in the 1%–5% range.
-Ang et al.
-(2023) find much larger optimal weights when investors have a preference for positive skewness, such as under constant relative risk aversion.
-Because the optimal crypto weight is sensitive to the joint return distribution—especially the correlation between cryptocurrencies and traditional assets—and because this correlation structure shifted after 2020 (see Fact 2), we re-estimate the theoretical optimal allocations using post-2020 data under the same Black–Litterman framework.
-Moreover, we extend the analysis to two representative investor classes: a retail investor and an institutional investor.
-In calculating the theoretical optimal crypto portfolio shares for them, we differ only in their investable options.
-This simplification abstracts from practical complications, such as liquidity, custody, and regulatory constraints, to make the exercise tractable.
-For a retail investor, we start with a baseline portfolio that is 50–50 between financial assets and real estate, with the financial portion 60–40 stocks–bonds.
-Based on the model, if we add an allocation to the cryptocurrency market, we would raise the portfolio’s Sharpe ratio from 0.39 to 0.52, with an optimal crypto weight of 3.1% of total wealth.
-As a historical counterfactual, we also compute the theoretical optimal portfolio using only pre-2020 data; the optimal crypto share is 1.4%.
-Mapping this to per-capita U.S.
-net wealth in 2019, a 1.4% allocation implies an investment of about \$3,500; held from 2020 to the end of our sample, this position would have grown by roughly \$52,000 (about 1,462%).
-This is a back-of-the-envelope, buy-and-hold calculation that assumes no rebalancing.
-For an institutional investor, we consider a broader investable universe—equities, government and corporate bonds, commodities, and real estate—with baseline weights proportional to each asset class’s market capitalization.
-Solving for the optimal portfolio and assuming slightly greater confidence in views on cryptocurrencies than in the retail case, the model implies an increase in the portfolio’s Sharpe ratio from 0.69 to 0.95, with an optimal share of 5.5% in the cryptocurrency market.
-These results show that—even with the post-2020 increase in comovement with equities—optimal allocations to cryptocurrency are sizable and in the 3–5% range, with larger weights for institutional portfolios.
-It is important to note that the estimations in this section are back-of-the-envelope and rely on the properties of returns that may not persist and abstract from important issues related to liquidity, custody, and regulatory risk, and therefore should be viewed as theoretical upper bounds on the potential diversification benefits of this emerging investable asset class, not as actual investment recommendations.
-For context, CalPERS, one of the largest U.S.
-pension funds, reports portfolio shares of comparable magnitude in emerging-market debt (5.3%), high-yield bonds (5.3%), private debt (3.8%), and mortgage-backed securities (5.8%) as of June 2025, indicating that the model-implied weights of cryptocurrencies are within a realistic range.
-3.4 Fact 4: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
-Fact 4: A small number of smart beta strategies—based on crypto-size, crypto-momentum, and crypto-value—span the cross-section of cryptocurrency returns.
+3.2 Stylized fact 2: How to be “smart” in crypto—crypto-size, crypto-momentum and crypto-value
+Stylized fact 2: A small number of smart beta strategies—based on crypto-size, crypto-momentum, and crypto-value—span the cross-section of cryptocurrency returns.
 A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors.
-In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French (1992))—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies.
-For cryptocurrencies, Liu et al.
-(2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns.
-Liu et al.
-(2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power.
+In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French 1992)—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies.
+For cryptocurrencies, Liu, Tsyvinski & Wu (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns.
+Liu, Tsyvinski & Wu (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power.
 We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
 We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE).
-[^3] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi et al.
-2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al.
+[^9] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi, Babiak & Dickerson 2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al.
 2022b; Borri & Santucci de Magistris 2022).
@@
 All three predictors generate statistically significant long-short returns in the full sample.
-The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu et al.
-(2022).
+The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu, Tsyvinski & Wu (2022).
 Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio.
@@
 The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns.
-For example, Asness et al.
-(2013) present evidence of value and momentum in equities, bonds, currencies, and commodities.
-The C-4 factor model of Liu et al.
-(2022) and Liu et al.
-(2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities.
+For example, Asness, Moskowitz & Pedersen (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities.
+The C-4 factor model of Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities.
 It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets.
 CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities).
-In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu et al.
-(2021) and proxy CVALUE with the price-to-new-address ratio.
-[^4] Liu et al.
-(2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
-Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong et al.
-2021b, Sockin & Xiong 2023a, b).
-A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”.
-Specifically, Liu et al.
-(2022) and Liu et al.
-(2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng et al.
-2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model.
-In Table 3, we study long-short strategies based on these signals post–2020.
-These strategies include portfolios based on long-horizon momentum, volatility, volume, and beta.
-We find that these long-short strategies continue not to generate significant average returns in the recent sample.
-[^5] In a recent study, Fieberg et al.
-(2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market.
-They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not.
-These results give evidence that the cross-section of cryptocurrency returns can be summarized by a small number of factors.
+In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu, Tsyvinski & Wu (2021) and proxy CVALUE with the price-to-new-address ratio.
+[^10] Liu, Tsyvinski & Wu (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
+Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong, Li & Wang 2021b; Sockin & Xiong 2023a, b).
+More broadly, this stylized low-dimensional structure is consistent with the broader multifactor view of asset pricing developed for equities, but the relevant priced risks in cryptocurrency include distinct, crypto-specific components—including network/adoption-related fundamentals.
 Table 2: Cryptocurrency smart beta strategies
@@
 *, **, and *** denote significance at the 10%, 5% and 1% levels.
+A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”.
+Specifically, Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng, Giglio & Xiu 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model.
+In Table 3, we study long-short strategies based on these signals post-2020.
+We find that these strategies continue not to generate significant average returns in the recent sample.
+[^11] In a recent study, Fieberg et al.
+(2024) examine 20,736 research designs for 43 sorting variables in the cryptocurrency market.
+They find that the most prominent cryptocurrency factors, such as size and momentum, remain consistently robust across various specifications, but the others do not.
+These results show that the cross-section of cryptocurrency returns can be summarized by a small number of factors.
 Table 3: Cryptocurrency smart beta strategies that are not beta
@@
 Notes: This table reports the mean quintile portfolio returns and long-short portfolio return for crypto smart beta strategies based on insignificant predictors.
-The predictors are past 12-week and 24-week return (MOM12 and MOM24), past 24-week return standard deviation (VOL), past 24-week dollar trading volume (VOLUME), coin market beta (CMKT BET) and stock market beta (MKT BETA) estimated using past 24-week returns.
+The predictors are past 12-week and 24-week return (MOM12 and MOM24), past 24-week return standard deviation (VOL), past 24-week dollar trading volume (VOLUME), coin market beta (CMKT BETA) and stock market beta (MKT BETA) estimated using past 24-week returns.
 The long-short portfolio is long portfolio 5 and short portfolio 1 for all predictors.
@@
 *, **, and *** denote significance at the 10%, 5% and 1% levels.
-3.5 Fact 5: Mind the Jumps—large, sudden price moves and “common disasters”
-Fact 5: Jumps are frequent and account for a large share of variation of cryptocurrency returns, far more than in equities. Thus, disasters are not rare but common.
+3.3 Stylized fact 3: Mind the Jumps—few factors, higher orders: why less is more
+Stylized fact 3: Cryptocurrency returns feature frequent jumps; in the cross-section, a parsimonious model augmented with a small number of higher-order terms captures most of the predictive content of black-box machine learning.
 Leibniz wrote that natura non facit saltus—nature makes no jumps.
 In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities.
-Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973), in their celebrated option pricing paper, to model prices as Brownian motions.
-[^6] Merton (1976) is aware of potential jumps in equity prices and include the feature in his paper.
-Rare disasters are also used as a central ingredient to explain risk-premia (e.g., Rietz (1988); Barro (2006)).
-Cryptocurrencies are different: jumps, and hence disasters, are not “rare disasters” but are “common disasters.” Specifically, Borri & Santucci de Magistris (2022) show that 12.8% of Bitcoin’s daily variability can be attributed to jumps—large, sudden price moves—and that higher-order moments such as skewness and kurtosis explain the risk premium on Bitcoin.
-[^7] See early evidence from Scaillet et al.
-(2020) similarly documenting very frequent jumps in Bitcoin.
-Fact 5 revisits and extend this finding, documenting the frequency and size of jumps in cryptocurrency and showing that those are far more relevant than for the U.S.
-stock market.
-Table 4 reports the frequency and magnitude of jumps in Bitcoin and Ethereum—the two largest cryptocurrencies by market capitalization—compared with U.S.
-stocks.
-Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006), a jump-robust measure of integrated variance.
-The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency.
-For U.S.
-equities, the same procedure is applied, and values represent cross-sectional averages across the constituents of the Dow Jones Industrial Average.
-Only jumps that are statistically significant at the 0.1% confidence level according to the BNS test of Barndorff-Nielsen & Shephard (2006) are retained.
-For Bitcoin and Ethereum, the probability of a significant jump on a given day is around 40 percent in the pre–2020 sample and 15–20 percent in the post–2020 period.
-In sharp contrast, the probability of a daily jump in U.S.
-stocks is below 10 percent before 2020 and falls to about 3 percent thereafter.
-Thus, jumps are far more frequent in cryptocurrencies than in equities.
-Conditional on a jump occurring, however, the share of total return variation explained by the jump is broadly similar across asset classes, particularly in the post–2020 period.
-This shows that what sets cryptocurrencies apart is the high frequency of jumps—contrary to Leibniz’s dictum that natura non facit saltus, cryptocurrencies are markets where saltus sunt regula (jumps are the rule), rather than the exception.
-These results show that higher-order moments are central to understanding cryptocurrency risk.
-Equity and bond markets follow relatively smooth dynamics with only rare discontinuities.
-Cryptocurrency markets are characterized by frequent and sizable jumps that are a defining feature of their return distribution.
-This prevalence of jumps, and the corresponding role of higher-order moments, means significant challenges for risk management and portfolio allocation (see, e.g., Borri 2019).
-The implication is that we need empirical asset-pricing strategies that explicitly accommodate higher-order moments and nonlinearities (see, e.g., Kozak et al.
-2020, Chetverikov et al.
-Forthcoming, Borri et al.
-2025).
+Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973) to model prices as Brownian motions.
+Cryptocurrencies are different: large, sudden price moves are common.
+Borri & Santucci de Magistris (2022) show that jumps account for a sizable share of Bitcoin’s return variation and that higher-order moments help explain its risk premium.
+[^12] See early evidence from Scaillet, Treccani & Trevisan (2020) similarly documenting very frequent jumps in Bitcoin.
+Table 4 reports the frequency and magnitude of statistically significant jumps in Bitcoin and Ethereum—relative to U.S.
+equities—using realized-variance and bipower-variation measures constructed from 5-minute data and the BNS test.
+Jumps occur far more frequently in cryptocurrencies than in the Dow Jones stocks, although conditional on a jump the share of variation attributable to the jump is broadly similar across assets.
+In short, contrary to Leibniz’s dictum, in crypto markets saltus sunt regula (jumps are the rule), not the exception.
 Table 4: Jumps in Cryptocurrency
@@
 stocks.
+Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006).
+The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency.
 Significant jumps are identified using the BNS test with a 0.1% significance level.
@@
 The post-2020 sample is the sample since Jan 1, 2020.
-3.6 Fact 6: Few factors, higher orders—rather than machine learning: why less is more
-Fact 6: In cryptocurrency, higher-order terms of the C-4 model accounts for approximately 25% of the cross-section of returns and captures most of the insights of machine learning methods.
-A growing literature in finance uses machine learning methods, such as neural networks, to investigate the cross-section of returns (see, for example, Gu et al.
-2020).
-Neural networks have become extremely successful in many prediction problems and moreover have the universal approximation property (Cybenko 1989; Hornik et al.
-1989)—a sufficiently large neural network can approximate any continuous function—and thus, in principle, capture any features of the data.
-These ideas have been present in the machine learning literature for several decades, but only recently have they become dramatically successful, as computational advances enabled the estimation of very large models, essentially aligning with the core idea of the universal approximation theorem.
-This has led finance and economics, as well as many other fields, to develop models with a very large number of parameters.
-The main shortcoming of neural networks for finance is that they are “black box” models.
-Once millions and sometimes hundred of millions of “factors” are run through often rather complicated and delicate gradient descent-like algorithms, it is difficult to see what features are economically important, how exactly the model produces results, and if an environment changes, whether the model still continues to perform.
-[^8] For a review of machine learning in asset pricing, see the book of Nagel (2021).
-One alternative to neural nets is the Kolmogorov–Arnold representation theorem (Kolmogorov 1956; Arnold 1957) introduced in finance by Borri et al.
+This prevalence of jumps, and the corresponding role of higher-order moments, poses challenges for risk management and portfolio choice (e.g., Borri 2019) and has a direct implication for asset pricing: nonlinearities matter.
+Motivated by this, a growing literature applies machine-learning methods—especially neural networks—to the cross-section of returns (e.g., Gu, Kelly & Xiu 2020).
+While neural networks have a universal approximation property (Cybenko 1989; Hornik, Stinchcombe & White 1989), their main drawback in finance is that they are “black box” models: they can fit complex patterns but make it difficult to isolate which economic features drive predicted returns and whether those relationships are stable when the environment changes.
+[^13] For a review of machine learning in asset pricing, see Nagel (2021).
+One alternative to neural nets is the Kolmogorov-Arnold (KA) representation theorem (Kolmogorov 1956; Arnold 1957), introduced in finance by Borri et al.
 (2024).
@@
 (2024), arXiv, April 11, 2024.
-Kolmogorov–Arnold remarkably shows that any nonlinear function can be written exactly as a combination of a finite sum of univariate functions composed with a link function.
-This formulation is thus a representation, not an approximation.
-Importantly, the early machine learning literature showed how to use in theory this representation to construct a shallower and smaller network (Hecht-Nielsen 1987, Girosi & Poggio 1989, Kurkova 1991).
-This network can be thought of as a much lower dimensional neural network where the activation functions of each neuron are being trained and adapt to the data.
-While theoretically this result is remarkable, [^10] In fact, Kolmogorov and Arnold construction solved Hilbert’s thirteenth problem, one of the major open problems of mathematics of the twentieth century.
-it is not clear whether this approach would work on actual data because of the complexity that the theory requires for the representation of any function, and in particular, on finance data, which is both relatively small and has a very low signal-to-noise ratio.
-Borri et al.
-(2024) is the first paper that shows the practical implementation and effectiveness of the Kolmogorov–Arnold architecture on actual data.
-Specifically, they show that this methodology enables construction of a single non-linear factor (KA factor) that captures a significant part of the variation in expected returns.
-[^11] Schmidt-Hieber (2021) is the first paper in the literature that studies quantitative properties of Kolmogorov–Arnold representation for multiple-layer neural networks and thus introduces the Kolmogorov–Arnold architecture.
-An early paper by Horowitz & Mammen (2007) discusses essentially a similar structure of general nonparametric regressions with an unknown link function.
+The theorem shows that a broad class of nonlinear functions can be exactly represented as finite sums of univariate functions composed with a link function—a structure that motivates shallow, low-dimensional architectures (see, e.g., Hecht-Nielsen 1987; Girosi & Poggio 1989; Kurkova 1991).
+In practice, Borri et al.
+(2024) implement this idea and construct a single nonlinear factor (the KA factor) that captures a sizable share of the cross-sectional variation in expected returns.
+[^15] Related work includes Schmidt-Hieber (2021) and Horowitz & Mammen (2007).
 Subsequent to Borri et al.
 (2024), Liu et al.
-(2024) (arXiv, April 30, 2024) popularized Kolmogorov Arnold Networks (KAN) to the machine learning community and showed how significantly shallower neural nets built on Kolmogorov–Arnold architecture may, in some cases, outperform and bring interpretability in several science applications.
-Toscano et al.
-(2025) argue that KAN rely on computation expensive basis functions with a performance that degrades with high parameter counts, which might limit the use in these science applications.
-The main advantage of this approach is that the KA-factor can be thought of as a simple representation of a much more complicated and larger neural network and, in principle, captures all such networks, without having to consider a variety of neural networks specifications and the optimization algorithms.
+(2024) popularize KAN in the machine-learning community, while Toscano, Wang & Karniadakis (2025) discuss computational and scaling limitations.
 Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion.
-However, the Taylor approximation to the higher order generates a very large set of factors, due to the need to account for all of the interactions of the higher-order terms in the Taylor expansion.
+However, the Taylor approximation to the higher order generates a very large set of factors, including all of the interactions of the higher-order terms in the Taylor expansion.
 Borri et al.
-(2025) develop the properties of a forward-selection Fama–MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns.
-The method proceeds sequentially, adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared.
-The main advantage of this approach is that it is a fully “glass box” model both in terms of determining the important factors and in the process to select them that allows to map features of the data to the economic factors and their higher order interactions.
+(2025) develop the properties of a forward-selection Fama-MacBeth (FS-FMB) method for selecting the higher-order factors and their interactions that are most important for the cross-section of returns by sequentially adding one factor or interaction to the selected set at a time and, at each step, choosing the term that maximizes the Fama–MacBeth (FMB) second-step cross-sectional R-squared.
+This approach is a fully “glass box” model in terms of determining the important factors.
 [^16] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian & Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017).
@@
 (2024) for an alternative “glass box” machine learning structure.
-Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu et al.
-(2021).
-The test assets are the set of 50 portfolios formed using the smart beta strategies described in Fact 4 (see Tables 2 and 3), and the CARB strategy described in Fact 8 (see Table 6, Panel B).
+Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu, Tsyvinski & Wu (2021).
+The test assets are the set of 50 portfolios formed using the smart beta strategies described in stylized fact 2 (see Tables 2 and 3), and the CARB strategy described in stylized fact 5 (see Table 6, Panel B).
 As with the classic CAPM, the C-1 model cannot account for the cross-section of average returns.
@@
 The procedure selects three higher-order factors.
-The first is CMOM cubed, resulting in an increase of the cross-sectional R-squared of approximately 15 pp.
-The importance of this factor is that it relates to the frequent and large jumps which are the distinct feature of cryptocurrency of Fact 5.
-The second and third selected factors are CMKT squared and the interaction between CMKT and CSIZE.
-Borri et al.
-(2025) show that the equity counterparts of these factors are important for the cross-section of equity returns.
 The cross-sectional R-squared of the FS-FMB procedure is 70%, a value very similar to that of the model including the KA-factor.
 In all models, the estimate of the intercept is not statistically different from zero.
-These results show that the FS–FMB procedure recovers essentially the same nonlinear structure as the KA single factor.
+These results show that the FS-FMB procedure recovers essentially the same nonlinear structure as the KA single factor.
 Moreover, since KA can represent any neural net, FS–FMB provides a simple, transparent, “open box” alternative to black-box models that is directly connected to the economic structure of the model and to observables characteristics.
@@
 This table reports the results from the second step of the FMB method.
-Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu et al.
-(2022).
+Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu, Tsyvinski & Wu (2022).
 The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic.
-The t-statistics is based on Newey-West corrected standard errors.
+The t-statistics are based on Newey-West corrected standard errors.
 Panel B refers to the C-4 model augmented by the single KA-factor, estimated with the KA procedure of Borri et al.
@@
 The method stops when the improvement in the R-squared in a given step is smaller than 1 pp.
-3.7 Fact 7: In crypto, the (block)chain drives the gain
-Fact 6: Blockchain economics and information significantly affects cryptocurrency prices, on-chain user adoption accounts for about 8% of the variation in cryptocurrency returns.
+3.4 Stylized fact 4: In crypto, the (block)chain drives the gain
+Stylized fact 4: Blockchain economics and information significantly affect cryptocurrency prices, on-chain user adoption accounts for about 8% of the variation in cryptocurrency returns.
 Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption.
 Classic work has analyzed how networks form and evolve (Jackson & Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet & Tirole, 2003), and how shocks propagate through financial networks (Allen & Gale, 2000).
-A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010, Acemoglu et al.
+A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010; Acemoglu et al.
 2011).
@@
 As Metcalfe’s law put it, “the value of a network grows as the square of the number of its users.”
-Empirically, Liu et al.
-(2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices.
+Empirically, Liu, Tsyvinski & Wu (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices.
 Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns.
-In time-series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
-Liu et al.
-(2021) also examine how markets absorb this information and what conditions amplify its relevance.
-Event-time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre-announcement behavior and no post-announcement drift—in contrast to equities.
+In time‑series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
+Liu, Tsyvinski & Wu (2021) also examine how markets absorb this information and what conditions amplify its relevance.
+Event‑time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre‑announcement behavior and no post‑announcement drift—in contrast to equities.
 The result gives evidence that investors fully incorporate the information at the weekly frequency.
-Lastly, Liu et al. (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in Fact 4.
-3.8 Fact 8: Young cryptocurrency markets, old inefficiencies.
-Fact 8: The young cryptocurrency markets are still inefficient.
+Lastly, Liu, Tsyvinski & Wu (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in stylized fact 2.
+Taken together, these results connect to classic theories of networks and information diffusion and suggest that network-based adoption and propagation forces are quantitatively important for cryptocurrency returns—arguably more so than in mature equity markets.
+3.5 Stylized fact 5: Young cryptocurrency markets, old inefficiencies.
+Stylized fact 5: The young cryptocurrency markets are still inefficient.
 But the origins are canonical: market segmentation and restrictions on capital flows.
@@
 A central concept in finance is the law of one price—identical securities should sell for identical prices.
-The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978, Harrison & Kreps 1979).
+The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978; Harrison & Kreps 1979).
 Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont & Thaler 2003), limits to arbitrage and market segmentation (Shleifer & Vishny 1997), or previously overlooked risks (Cochrane 2002).
@@
 These deviations—classic examples of market inefficiency and violations of the law of one price—are results of market segmentation induced by capital controls and idiosyncratic domestic shocks that prevent arbitrage.
+Fragmentation and limits to arbitrage also have implications for market fragility and flash-crash dynamics.
+Menkveld & Yueshen (2019) study the May 6, 2010 U.S.
+equity Flash Crash and argue that a disruption in cross-market arbitrage can make fragmented markets fragile and result in price crashes; they also refer to analogous flash-crash episodes in cryptocurrency markets, including in Ethereum (June 21, 2017) and Bitcoin (October 10, 2017).
+Because our baseline analysis is conducted at the weekly frequency using close-to-close returns, these very short-lived intraday dislocations are less likely to affect our main results, though they remain relevant for studies of market microstructure and tail risk.
 [^17] Barbon & Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
@@
 Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly.
-Specifically, Borri & Shakhnov (2022) shows that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe.
+Specifically, Borri & Shakhnov (2022) show that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe.
 That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer & Vishny (1997).
 As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades.
-[^14] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman Fried.
+[^18] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman-Fried.
 However, Borri & Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
-We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish łoty, Swiss franc, British pound, and U.S.
+We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish złoty, Swiss franc, British pound, and U.S.
 dollar).
@@
 Borri & Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
-The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu et al.
-(2023), is the most visible manifestation of the results described in this fact.
+The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu, Makarov & Schoar (2023), is the most visible manifestation of the results described in this fact.
+More broadly, the evidence from cryptocurrency markets provides a clean illustration of the classic limits-to-arbitrage view: when trading, settlement, and capital-mobility frictions are economically large, violations of the law of one price can persist rather than being immediately arbitraged away.
 Table 6: Cryptocurrency “Arbitrage” Strategy
@@
 *, **, and *** denote significance at the 10%, 5% and 1% levels.
-3.9 Fact 9: When the funding dries up, we finally learn the worth of futures
-Fact 9: Cryptocurrency carry trade, the defining characteristics of the futures market, used to be highly profitable, but profitability has compressed sharply since 2024.
-Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973, Merton 1973, Black 1976).
+3.6 Stylized fact 6: When the funding dries up, we finally learn the worth of futures
+Stylized fact 6: Cryptocurrency carry trade, the defining characteristics of the futures market, used to be highly profitable, but profitability has compressed sharply since 2024.
+Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973; Merton 1973; Black 1976).
 There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs.
-Among all derivatives, perpetual futures are the most common derivative contract in cryptocurrency markets.
-For example, Schmeling et al.
-(2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample.
+Among all derivatives, perpetual futures are the most common derivative contracts in cryptocurrency markets.
+For example, Schmeling, Schrimpf & Todorov (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample.
 Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot.
-Ackerer et al.
-(2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures.
-The crypto-carry trade strategy of Schmeling et al.
-(2023) is short a perpetual futures contract and long a position in the corresponding spot market.
+Ackerer, Hugonnier & Jermann (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures.
+The crypto-carry trade strategy of Schmeling, Schrimpf & Todorov (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
 Figure 2 plots the evolution over time of one dollar invested in the crypto-carry strategy (black line, left y-axis) and in the spot Bitcoin market (red line, right y-axis).
@@
 The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis).
-The Carry Trade is defined as in Schmeling et al.
-(2023).
+The Carry Trade is defined as in Schmeling, Schrimpf & Todorov (2023).
 The sample goes from August 1, 2020 to May 31, 2025.
-3.10 Fact 10: Growing up with supervision—regulation and oversight strengthen cryptocurrency markets
-Fact 10: As the market matures, there is growing disclosure and transparency in cryptocurrencies.
+3.7 Stylized fact 7: Growing up with supervision—regulation and oversight strengthen cryptocurrency markets
+Stylized fact 7: As the market matures, there is growing disclosure and transparency in cryptocurrencies.
 A fundamental insight of information economics is that markets with asymmetric information can fail to allocate resources efficiently.
 When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970).
-Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983, Dye 1985, Healy & Palepu 2001, Verrecchia 2001).
-However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001, Ma & Shleifer 2025).
+Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983; Dye 1985; Healy & Palepu 2001; Verrecchia 2001).
+However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001; Ma & Shleifer 2025).
 Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information.
@@
 2025).
-On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021, Liu et al.
-2025, Du et al.
-2025).
-Howell et al.
-(2020) and Bourveau et al.
+On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021; Liu, Sheng & Wang 2025; Du, Huang & Wu 2025).
+Howell, Niessner & Yermack (2020) and Bourveau et al.
 (2022) show the importance of voluntary disclosures in the initial coin offering stage.
 Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al.
-2022, Bourveau et al.
-2024).
+2022; Bourveau, Brendel & Schoenfeld 2024).
 In July 2025, Congress enacted the Guiding and Establishing National Innovation for U.S.
@@
 4 ADDITIONAL DISCUSSIONS
-The ten facts summarize key developments in cryptocurrencies through the lens of empirical asset pricing.
+The seven stylized facts summarize key developments in cryptocurrencies through the lens of empirical asset pricing.
 In this section, we provide additional discussions on emerging topics, including retail investor trading in the cryptocurrency market, the non-fungible token market, and studies on the dark side of cryptocurrency usage.
@@
 households, showing that cryptocurrency investors differ systematically from other investors.
-Alvarez et al.
-(2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption.
-Han et al.
-(2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets.
+Alvarez, Argente & Van Patten (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption.
+Han, Liu & Sui (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets.
 Finally, Benetton & Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies.
@@
 A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs).
-Goetzmann & Nozari (2022) is the first to construct repeated sales return indices of NFTs and examine their properties.
+Goetzmann & Nozari (2022) are the first to construct repeated sales return indices of NFTs and examine their properties.
 Borri et al.
@@
 Furthermore, Borri et al.
-(2022a) shows that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market.
+(2022a) show that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market.
 Huang & Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing.
-Oh et al.
-(2025) document preference-driven herding in NFT primary market sales.
+Oh, Rosen & Zhang (2025) document preference-driven herding in NFT primary market sales.
 A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years.
-Foley et al.
-(2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley et al.
-(2019) are largely driven by double-counting.
+Foley, Karlsen & Putniņš (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley, Karlsen & Putniņš (2019) are largely driven by double-counting.
 Griffin & Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation.
-Amiram et al.
-(2022) show that abnormal spikes in on-chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, and Cong et al.
-(2025) provide an anatomy of crypto-enabled crimes.
-Li et al.
-(2025) document pervasive cryptocurrency pump-and-dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre-event run-ups consistent with insider gains at outsiders’ expense.
+Amiram, Jørgensen & Rabetti (2022) show that abnormal spikes in on‑chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, Cong et al.
+(2025) provide an anatomy of crypto-enabled crimes and identify ransomware as the dominant activity, while Cong et al.
+(2023a) review crypto-related scams, including Ponzi schemes and “rug pulls” in decentralized projects and token offerings.
+Li, Shin & Wang (2025) document pervasive cryptocurrency pump‑and‑dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre‑event run‑ups consistent with insider gains at outsiders’ expense.
 5 CONCLUSIONS
-In the 16 years since Nakamoto, cryptocurrencies have reached a certain stage of maturity: they can now be studied using the same measurement, data quality, and research rigor applied to other asset classes.
-The market is now worth more than four trillion dollars.
+In the 17 years since Nakamoto, cryptocurrencies have reached a certain stage of maturity: they can now be studied using the same measurement, data quality, and research rigor applied to other asset classes.
+By the end of our sample period (September 2025), aggregate cryptocurrency market capitalization is about \$4 trillion.
 Major financial firms are using cryptocurrency in their businesses.
-We show that there are important similarities between cryptocurrency markets and traditional asset markets—Sharpe ratios are broadly comparable, and a small number of factors and their higher orders summarize the cross-section of returns.
+We show that there are important similarities between cryptocurrency markets and traditional asset markets—Sharpe ratios so far are broadly comparable, and a small number of factors and their higher orders summarize the cross-section of returns.
 On the other hand, cryptocurrency also has its own character—frequent jumps and blockchain information driving prices.
@@
 The right way to understand it is exactly through traditional finance tools—empirical asset pricing.
-The ten facts we present summarize the key takeaways from this view.
+The seven stylized facts we present summarize the key takeaways from this view.
 DISCLOSURE STATEMENT

### v4

@@
 Money has long been a central topic in economics—for example, see classic contributions of Tobin (1965) and Friedman (1968) to more recent advances.
-A large literature studied why money exists and how it functions—the role of money in economies with spatially separated agents (Townsend 1980), the coexistence of different types of money under different institutional restrictions (Wallace 1983), search-based models (Kiyotaki & Wright, 1989; Trejos & Wright, 1995; Shi, 1995; Lagos & Wright, 2005), mechanism-design approaches (Green & Zhou, 1998), the “money-as-memory” view (Kocherlakota, 1998), commodity and token money (Sargent & Wallace, 1983; Sargent, 2019), competing currencies (Kareken & Wallace, 1981), and the money-like safe assets (Gorton & Pennacchi 1990; Holmström & Tirole 1998).
+A large literature studied why money exists and how it functions—the role of money in economies with spatially separated agents (Townsend 1980), the coexistence of different types of money under different institutional restrictions (Wallace 1983), search-based models (Kiyotaki and Wright, 1989; Trejos and Wright, 1995; Shi, 1995; Lagos and Wright, 2005), mechanism-design approaches (Green and Zhou, 1998), the “money-as-memory” view (Kocherlakota, 1998), commodity and token money (Sargent and Wallace, 1983; Sargent, 2019), competing currencies (Kareken and Wallace, 1981), and the money-like safe assets (Gorton and Pennacchi 1990; Holmström and Tirole 1998).
 When the paper of Nakamoto (2008) introduced a permissionless blockchain that provides public record-keeping without a central authority, financed and governed by a native token, the canonical literature on money had reached a degree of maturity.
@@
 When practice advances faster than theory, it is natural to start asking empirical questions that can themselves later be used to shape theory.
-The paper of Liu & Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing.
-A sizable literature uses an asset pricing lens to study risk exposures, cross‑sectional return patterns, cash‑flow‑like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov & Schoar 2020; Liu, Tsyvinski & Wu 2022; Alvarez, Argente & Van Patten 2023; Sockin & Xiong 2023a; von Luckner, Reinhart & Rogoff 2023; Kogan et al.
+The paper of Liu and Tsyvinski (2021) is the first comprehensive analysis of cryptocurrency through the lens of empirical finance and asset pricing.
+A sizable literature uses an asset pricing lens to study risk exposures, cross‑sectional return patterns, cash‑flow‑like primitives, market microstructure, and limits to arbitrage (see, for example, Makarov and Schoar 2020; Liu, Tsyvinski, and Wu 2022; Alvarez, Argente, and Van Patten 2023; Sockin and Xiong 2023b; von Luckner, Reinhart, and Rogoff 2023; Kogan et al.
 2024).
@@
 When cryptocurrency first appeared it was easy to dismiss it: money engineered by an unknown author and operating outside of the traditional financial system.
-[^1] See, Foley, Karlsen & Putniņš (2019) and Griffin & Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations.
+[^1] See, Foley, Karlsen, and Putniņš (2019) and Griffin and Shams (2020) for early analyses documenting how cryptocurrency mainly enabled illicit activity and alleged market manipulations.
 It is worth noting that several important financial innovations, such as the limited‑liability company and mortgage‑backed securities, were also initially met with distrust before becoming part of standard finance (see, for example, Hilt 2014 and Lewis 2010), as institutions and legal frameworks evolved to manage their risks and as research deepened understanding of these innovations.
-Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi & Brunnermeier 2018; Leshno & Strack 2020; Biais et al.
-2019; Kogan, Fanti & Viswanath 2022; Brunnermeier, James & Landau 2019; Benigno, Schilling & Uhlig 2022; Capponi, Jia & Wang 2025; Jermann & Xiang 2025), regulation (Cong & He 2019; Townsend 2020; Auer, Monnet & Shin 2025), digital asset pricing and valuation (Athey et al.
-2016; Cong, Li & Wang 2021b; Biais et al.
-2023a), transaction costs (Easley, O’Hara & Basu 2019; Cong, He & Li 2021a; Basu et al.
+Since then research has evolved toward a much broader agenda that increasingly resembles research on any other asset class—the market structure (Abadi and Brunnermeier 2018; Leshno and Strack 2020; Biais et al.
+2019; Kogan, Fanti, and Viswanath 2022; Brunnermeier, James, and Landau 2019; Benigno, Schilling, and Uhlig 2022; Capponi, Jia, and Wang 2025; Jermann and Xiang 2025), regulation (Cong and He 2019; Townsend 2020; Auer, Monnet, and Shin 2025), digital asset pricing and valuation (Athey et al.
+2016; Cong, Li, and Wang 2021b; Biais et al.
+2023), transaction costs (Easley, O’Hara, and Basu 2019; Cong, He, and Li 2021a; Basu et al.
 2023), central bank money (Auer et al.
-2022; Whited, Wu & Xiao 2023; Schilling, Fernández-Villaverde & Uhlig 2024), securitization (Cong, Li & Wang 2022; Sockin & Xiong 2023a; Rogoff & You 2023; Goldstein, Gupta & Sverchkov 2024), safe assets (Ma, Zeng & Zhang 2025), international capital flow (von Luckner, Reinhart & Rogoff 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde & Sanches 2018) and market microstructure (Easley et al.
+2022; Whited, Wu, and Xiao 2023; Schilling, Fernández-Villaverde, and Uhlig 2024), securitization (Cong, Li, and Wang 2022; Sockin and Xiong 2023b; Rogoff and You 2023; Goldstein, Gupta, and Sverchkov 2024), safe assets (Ma, Zeng, and Zhang 2025), international capital flow (von Luckner, Reinhart, and Rogoff 2023), blockchain technologies and the traditional financial system (Fernández-Villaverde and Sanches 2018) and market microstructure (Easley et al.
 2024).
@@
 These stylized facts help evaluate and revisit both classic and new theories of money, while remaining conditional on the available data and institutional environment to date and therefore potentially evolving as market structure, regulation, and adoption change.
-[^2] For complementary surveys on blockchain architecture, decentralized finance and central bank digital currencies (CBDC), see Kose et al.
-(2022); Makarov et al.
-(2022); Kose et al.
-(2023); Biais et al.
-(2023b); Adrian & Mancini-Griffoli (2021); Infante et al.
-(2023).
 Seven Stylized Facts in Cryptocurrency Asset Pricing and Empirical Finance.
@@
 The final dataset comprises 16,468 unique coins and approximately 9.5 million records.
-Second, for the construction of the crypto carry strategy of Schmeling, Schrimpf & Todorov (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025.
-The data on new addresses are from Liu, Tsyvinski & Wu (2021).
-Moreover, for the construction of the crypto arbitrage strategy of Borri & Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025.
+Second, for the construction of the crypto carry strategy of Schmeling, Schrimpf, and Todorov (2023), we collect perpetual futures prices, funding rates and spot index prices for Bitcoin at the 8-hour frequency from Binance for the period August 1, 2020 to May 31, 2025.
+The data on new addresses are from Liu, Tsyvinski, and Wu (2021).
+Moreover, for the construction of the crypto arbitrage strategy of Borri and Shakhnov (2022), we collect Bitcoin prices against 510 pairs from 165 different centralized exchanges quoted against 49 fiat currencies from CryptoCompare.com for the period December 31, 2013 to January 30, 2025.
 We convert these quotes into U.S.
 dollars using spot FX rates from Bloomberg.
-Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.333Finally, we obtain price indices for the MSCI global stock market, global high-yield and government bonds, a broad commodity index, gold, the U.S.
+Furthermore, for estimation of jumps in cryptocurrency we collect tick-level transaction data for Bitcoin and Ethereum from Kraken.com and measures of realized variance constructed from tick-level data for the components of the Dow Jones Industrial Average Index from the CaPiRe dataset, for the period January 1, 2015 to March 31, 2025.222Finally, we obtain price indices for the MSCI global stock market, global high-yield and government bonds, a broad commodity index, gold, the U.S.
 dollar index, the Dow Jones US Real Estate Index, the U.S.
@@
 A strand of the empirical cryptocurrency literature argues that parts of the historical crypto data—especially in early years—may reflect manipulation or unreliable reporting, including alleged price manipulation on Mt.
-Gox in 2013 (Gandal et al., 2018), inflated or “artificial” trading volume reported by some exchanges (Makarov & Schoar, 2020), and exchange-managed wash trading concentrated on “unregulated” exchanges and largely absent on “regulated” exchanges such as Bitstamp, Coinbase, and Gemini (Cong et al., 2023b), as well as trade reordering/front-running on decentralized exchanges (Daian et al., 2019).
-At the same time, on-chain evidence points to substantial cross-exchange flows: analyzing the Bitcoin blockchain, Makarov & Schoar (2021) show that a sizable share of volume to exchanges originates from other exchanges or large market participants, consistent with arbitrage activity that tends to limit persistent price deviations across major venues.
+Gox in 2013 (Gandal et al., 2018), inflated or “artificial” trading volume reported by some exchanges (Makarov and Schoar, 2020), and exchange-managed wash trading concentrated on “unregulated” exchanges and largely absent on “regulated” exchanges such as Bitstamp, Coinbase, and Gemini (Cong et al., 2023b), as well as trade reordering/front-running on decentralized exchanges (Daian et al., 2019).
+At the same time, on-chain evidence points to substantial cross-exchange flows: analyzing the Bitcoin blockchain, Makarov and Schoar (2021) show that a sizable share of volume to exchanges originates from other exchanges or large market participants, consistent with arbitrage activity that tends to limit persistent price deviations across major venues.
 While these concerns cannot be fully eliminated, we mitigate them in several ways.
@@
 Gox manipulation episode in 2013.
-Second, our pre-/post-2020 split places emphasis on the most recent period, and the post-2020 subsample begins after the main evidence documented in Makarov & Schoar (2020) and Cong et al.
+Second, our pre-/post-2020 split places emphasis on the most recent period, and the post-2020 subsample begins after the main evidence documented in Makarov and Schoar (2020) and Cong et al.
 (2023b).
@@
 From the classic finance point of view, the relevant measure is, however, risk-adjusted performance—the Sharpe ratio, i.e., mean excess return per unit of volatility (Sharpe 1966).
-Liu & Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable.
-Stylized fact 1 revisits and confirms this finding in a longer dataset through September 2025.777These performance measures are computed from price indices and thus reflect the return to the underlying asset, not the return to a particular investor after custody risk.
+Liu and Tsyvinski (2021) establish that the aggregate cryptocurrency market exhibits returns and volatilities an order of magnitude higher than those of the aggregate stock market, but Sharpe ratios remain broadly comparable.
+Stylized fact 1 revisits and confirms this finding in a longer dataset through September 2025.666These performance measures are computed from price indices and thus reflect the return to the underlying asset, not the return to a particular investor after custody risk.
 Episodes such as exchange cybersecurity breaches or bankruptcies (e.g., Mt.
@@
 Table 1, Panel B, reports correlations between the coin market return and major asset classes.
-In the pre-2020 period, correlations with traditional assets are small and statistically insignificant, consistent with Liu & Tsyvinski (2021).
+In the pre-2020 period, correlations with traditional assets are small and statistically insignificant, consistent with Liu and Tsyvinski (2021).
 In the post-2020 period, correlations rise sharply—most notably with global equities, increasing from 1% pre-2020 to 39% post-2020—and the coin market also becomes significantly correlated with commodities (27%), high-yield bonds (24%), government bonds (13%), and gold (18%).
@@
 Since this measure incorporates an inflation risk premium, the result may reflect the same underlying risk component that drives the stronger comovement between cryptocurrencies and equities in recent years.
-The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong, Fang & Lin (2023) and Jiang et al.
+The increase in correlation between the cryptocurrency market and traditional asset classes has been documented in recent studies, for example, Dong, Fang, and Lin (2023) and Jiang et al.
 (2025).
@@
 In mean–variance settings, an asset’s contribution to diversification depends on its covariance with the rest of the portfolio rather than its standalone volatility (Markowitz, 1952).
-Consistent with the portfolio results in Liu & Tsyvinski (2021) and related Bayesian approaches (Duchin et al., 2025), a simple Black–Litterman calibration using post-2020 moments still implies non-trivial but modest allocations to the aggregate crypto market—about 3.1% for a representative retail portfolio and 5.5% for an institutional portfolio with a broader opportunity set.
+Consistent with the portfolio results in Liu and Tsyvinski (2021) and related Bayesian approaches (Duchin et al., 2025), a simple Black–Litterman calibration using post-2020 moments still implies non-trivial but modest allocations to the aggregate crypto market—about 3.1% for a representative retail portfolio and 5.5% for an institutional portfolio with a broader opportunity set.
 We emphasize that these magnitudes are back-of-the-envelope and depend on strong assumptions about expected returns, correlations, and investability (liquidity, custody, and regulatory constraints), and should therefore be interpreted as illustrative upper bounds on diversification benefits rather than investment advice.
@@
 A central question in asset pricing is whether systematic variation in expected returns can be explained by a parsimonious set of factors.
-In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama & French 1992)—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies.
-For cryptocurrencies, Liu, Tsyvinski & Wu (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns.
-Liu, Tsyvinski & Wu (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power.
+In equities, substantial academic endeavor—most prominently by Eugene Fama and Kenneth French (e.g., Fama and French 1992)—established this approach, and factor models play a core role in modern finance practice and theory from risk management of large banks to niche hedge fund strategies.
+For cryptocurrencies, Liu, Tsyvinski, and Wu (2022) establish that a three factor model consisting of the aggregate cryptocurrency market, size, and momentum (C-3 factor model), captures a large fraction of the cross-section of expected returns.
+Liu, Tsyvinski, and Wu (2021) show that adding a value factor (C-4 factor model) further strengthens the explanatory power.
 We revisit this evidence using our updated sample and show that a small number of predictors drive the cross-section of cryptocurrency returns over the longer sample and post-2020.
 We construct smart beta strategies by sorting coins into quintile portfolios based on three established significant predictors: market capitalization (CSIZE), past two-week returns (CMOM), and the price-to-new-address ratio (CVALUE).
-[^9] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi, Babiak & Dickerson 2022), market segmentation (e.g., Borri & Shakhnov 2022), and higher-order moments (e.g., Borri et al.
-2022b; Borri & Santucci de Magistris 2022).
+[^8] Studies based on smaller samples of coins have also proposed additional predictors of cryptocurrency returns in the cross-section, including liquidity (e.g., Bianchi, Babiak, and Dickerson 2022), market segmentation (e.g., Borri and Shakhnov 2022), and higher-order moments (e.g., Borri et al.
+2022b; Borri and Santucci de Magistris 2022).
 The return to each predictor is measured as the long-short spread between the highest and lowest quintiles (Portfolio 5 minus Portfolio 1).
@@
 All three predictors generate statistically significant long-short returns in the full sample.
-The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu, Tsyvinski & Wu (2022).
+The average long-short returns based on size and value are significantly negative and the average long-short return based on momentum is significantly positive, consistent with Liu, Tsyvinski, and Wu (2022).
 Cryptocurrency value is the weekly quintile portfolio of each bracket based on the price-to-new address ratio.
@@
 The C-4 structure of the crypto-market thus parallels the equity market, where a small number of equity-market factors—most notably size, value, momentum, and profitability—dominate the cross-section of expected returns.
-For example, Asness, Moskowitz & Pedersen (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities.
-The C-4 factor model of Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama & French (1992) and Carhart (1997) for equities.
+For example, Asness, Moskowitz, and Pedersen (2013) present evidence of value and momentum in equities, bonds, currencies, and commodities.
+The C-4 factor model of Liu, Tsyvinski, and Wu (2022) and Liu, Tsyvinski, and Wu (2021) is the cryptocurrency counterpart of the classic 4-factor model of Fama and French (1992) and Carhart (1997) for equities.
 It is important to note that the predictors in cryptocurrency markets are only partly analogous to those in equity markets.
 CSIZE and CMOM closely resemble their equity counterparts, although crypto momentum is measured over a much shorter horizon (previous two weeks, versus the standard 12-to-2 months in equities).
-In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu, Tsyvinski & Wu (2021) and proxy CVALUE with the price-to-new-address ratio.
-[^10] Liu, Tsyvinski & Wu (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
-Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong, Li & Wang 2021b; Sockin & Xiong 2023a, b).
+In contrast, there is no direct analogue to book value (Shiller 1981) in cryptocurrencies, so we follow Liu, Tsyvinski, and Wu (2021) and proxy CVALUE with the price-to-new-address ratio.
+[^9] Liu, Tsyvinski, and Wu (2021) show that the number of new addresses captures the network effects of cryptocurrencies and serves as a measure of their fundamental value.
+Establishing the small factor structure of the cryptocurrency returns, in our view, is not only important from purely the empirical asset pricing point of view but also since it provides a set of stylized low dimensional features that capture the main insights of several important theoretical models of cryptocurrency (Cong, Li, and Wang 2021b; Sockin and Xiong 2023b, a).
 More broadly, this stylized low-dimensional structure is consistent with the broader multifactor view of asset pricing developed for equities, but the relevant priced risks in cryptocurrency include distinct, crypto-specific components—including network/adoption-related fundamentals.
@@
 A related important point is that in addition to determining which betas are smart, it is equally as important to establish which betas from the factor zoo are not “smart”.
-Specifically, Liu, Tsyvinski & Wu (2022) and Liu, Tsyvinski & Wu (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng, Giglio & Xiu 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model.
+Specifically, Liu, Tsyvinski, and Wu (2022) and Liu, Tsyvinski, and Wu (2021) show that—besides crypto-specific size, momentum, and value—most of the smart beta strategies from the canonical factor zoo in the equity market (Feng, Giglio, and Xiu 2020) either do not generate significant hedged long-short strategy returns or are subsumed by the C-4 factor model.
 In Table 3, we study long-short strategies based on these signals post-2020.
@@
 In economics and finance, this often means that most changes occur gradually, with only occasional discontinuities.
-Indeed, jumps are rare in many asset classes, and this key insight leads Black & Scholes (1973) to model prices as Brownian motions.
+Indeed, jumps are rare in many asset classes, and this key insight leads Black and Scholes (1973) to model prices as Brownian motions.
 Cryptocurrencies are different: large, sudden price moves are common.
-Borri & Santucci de Magistris (2022) show that jumps account for a sizable share of Bitcoin’s return variation and that higher-order moments help explain its risk premium.
-[^12] See early evidence from Scaillet, Treccani & Trevisan (2020) similarly documenting very frequent jumps in Bitcoin.
+Borri and Santucci de Magistris (2022) show that jumps account for a sizable share of Bitcoin’s return variation and that higher-order moments help explain its risk premium.
+[^11] See early evidence from Scaillet, Treccani, and Trevisan (2020) similarly documenting very frequent jumps in Bitcoin.
 Table 4 reports the frequency and magnitude of statistically significant jumps in Bitcoin and Ethereum—relative to U.S.
@@
 stocks.
-Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen & Shephard (2002) and the bi-power variation of Barndorff-Nielsen & Shephard (2006).
+Jumps are identified as the difference between the daily realized variance of Barndorff-Nielsen and Shephard (2002) and the bi-power variation of Barndorff-Nielsen and Shephard (2006).
 The two realized measures are constructed from tick-level data aggregated at the 5-minute frequency.
@@
 This prevalence of jumps, and the corresponding role of higher-order moments, poses challenges for risk management and portfolio choice (e.g., Borri 2019) and has a direct implication for asset pricing: nonlinearities matter.
-Motivated by this, a growing literature applies machine-learning methods—especially neural networks—to the cross-section of returns (e.g., Gu, Kelly & Xiu 2020).
-While neural networks have a universal approximation property (Cybenko 1989; Hornik, Stinchcombe & White 1989), their main drawback in finance is that they are “black box” models: they can fit complex patterns but make it difficult to isolate which economic features drive predicted returns and whether those relationships are stable when the environment changes.
+Motivated by this, a growing literature applies machine-learning methods—especially neural networks—to the cross-section of returns (e.g., Gu, Kelly, and Xiu 2020).
+While neural networks have a universal approximation property (Cybenko 1989; Hornik, Stinchcombe, and White 1989), their main drawback in finance is that they are “black box” models: they can fit complex patterns but make it difficult to isolate which economic features drive predicted returns and whether those relationships are stable when the environment changes.
 [^12] For a review of machine learning in asset pricing, see Nagel (2021).
@@
 (2024), arXiv, April 11, 2024.
-The theorem shows that a broad class of nonlinear functions can be exactly represented as finite sums of univariate functions composed with a link function—a structure that motivates shallow, low-dimensional architectures (see, e.g., Hecht-Nielsen 1987; Girosi & Poggio 1989; Kurkova 1991).
+The theorem shows that a broad class of nonlinear functions can be exactly represented as finite sums of univariate functions composed with a link function—a structure that motivates shallow, low-dimensional architectures (see, e.g., Hecht-Nielsen 1987; Girosi and Poggio 1989; Kurkova 1991).
 In practice, Borri et al.
 (2024) implement this idea and construct a single nonlinear factor (the KA factor) that captures a sizable share of the cross-sectional variation in expected returns.
-[^15] Related work includes Schmidt-Hieber (2021) and Horowitz & Mammen (2007).
+[^14] Related work includes Schmidt-Hieber (2021) and Horowitz and Mammen (2007).
 Subsequent to Borri et al.
 (2024), Liu et al.
-(2024) popularize KAN in the machine-learning community, while Toscano, Wang & Karniadakis (2025) discuss computational and scaling limitations.
+(2024) popularize KAN in the machine-learning community, while Toscano, Wang, and Karniadakis (2025) discuss computational and scaling limitations.
 Another alternative to neural networks follows the traditional finance toolbox—approximate and estimate the stochastic discount factor via a Taylor expansion.
@@
 This approach is a fully “glass box” model in terms of determining the important factors.
-[^16] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian & Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017).
+[^15] Conceptually, these higher-order interaction terms capture conditional moments, as in systemic-risk measures such as CoVaR (Adrian and Brunnermeier, 2016) and marginal expected shortfall (Acharya et al., 2017).
 See Bell et al.
 (2024) for an alternative “glass box” machine learning structure.
-Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu, Tsyvinski & Wu (2021).
+Table 5, Panel A, reports second-step FMB results for the crypto-CAPM, or C-1, model and for the C-4 model of Liu, Tsyvinski, and Wu (2021).
 The test assets are the set of 50 portfolios formed using the smart beta strategies described in stylized fact 2 (see Tables 2 and 3), and the CARB strategy described in stylized fact 5 (see Table 6, Panel B).
@@
 This table reports the results from the second step of the FMB method.
-Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu, Tsyvinski & Wu (2022).
+Panel A refers to baseline models: the crypto-CAPM (C-1) and the C-4 model of Liu, Tsyvinski, and Wu (2022).
 The table reports the adjusted cross-sectional R-squared, the estimate for the intercept ($\alpha$) and associated t-statistic.
@@
 Network externalities are a fundamental concept in economics because they shape markets and influence competition, regulation and technology adoption.
-Classic work has analyzed how networks form and evolve (Jackson & Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet & Tirole, 2003), and how shocks propagate through financial networks (Allen & Gale, 2000).
-A central question common in this work is information diffusion and learning in a network (Golub & Jackson 2010; Acemoglu et al.
+Classic work has analyzed how networks form and evolve (Jackson and Wolinsky, 1996), how pricing and competition operate on networked platforms (Rochet and Tirole, 2003), and how shocks propagate through financial networks (Allen and Gale, 2000).
+A central question common in this work is information diffusion and learning in a network (Golub and Jackson 2010; Acemoglu et al.
 2011).
@@
 As Metcalfe’s law put it, “the value of a network grows as the square of the number of its users.”
-Empirically, Liu, Tsyvinski & Wu (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices.
+Empirically, Liu, Tsyvinski, and Wu (2021) show that on-chain user adoption—measured as new user growth—has a statistically and economically significant effect on cryptocurrency prices.
 Importantly, this single on-chain variable accounts for about 8% of the variation in cryptocurrency returns.
 In time‑series tests for the largest coins, on-chain user adoption alone can account for more than 10% of the variation in their returns.
-Liu, Tsyvinski & Wu (2021) also examine how markets absorb this information and what conditions amplify its relevance.
+Liu, Tsyvinski, and Wu (2021) also examine how markets absorb this information and what conditions amplify its relevance.
 Event‑time evidence shows a sharp price reaction when weekly adoption information is realized, with muted pre‑announcement behavior and no post‑announcement drift—in contrast to equities.
 The result gives evidence that investors fully incorporate the information at the weekly frequency.
-Lastly, Liu, Tsyvinski & Wu (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in stylized fact 2.
+Lastly, Liu, Tsyvinski, and Wu (2021) show that the crypto value measure, constructed as the price-to-new-address ratio, negatively predicts cryptocurrency returns—a result we discuss in stylized fact 2.
 Taken together, these results connect to classic theories of networks and information diffusion and suggest that network-based adoption and propagation forces are quantitatively important for cryptocurrency returns—arguably more so than in mature equity markets.
@@
 A central concept in finance is the law of one price—identical securities should sell for identical prices.
-The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978; Harrison & Kreps 1979).
-Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont & Thaler 2003), limits to arbitrage and market segmentation (Shleifer & Vishny 1997), or previously overlooked risks (Cochrane 2002).
+The law of one price, and the related concept of absence of arbitrage, is important for the existence of a discount factor (Ross 1978; Harrison and Kreps 1979).
+Deviations from the law of one price have been presented as proof of irrational investor behavior (Lamont and Thaler 2003), limits to arbitrage and market segmentation (Shleifer and Vishny 1997), or previously overlooked risks (Cochrane 2002).
 One of the main features of cryptocurrency is that it is possible to trade coins simultaneously on multiple exchanges and against a number of fiat and crypto pairs.
@@
 dollar or the euro, or even against another cryptocurrency such as Ethereum or Tether, a stablecoin.
-Makarov & Schoar (2020) and Borri & Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs.
+Makarov and Schoar (2020) and Borri and Shakhnov (2023) document sizable, persistent deviations in the dollar price of Bitcoin across exchanges and across fiat- and crypto-quoted pairs.
 We revisit this evidence on updated data for 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies.
@@
 Fragmentation and limits to arbitrage also have implications for market fragility and flash-crash dynamics.
-Menkveld & Yueshen (2019) study the May 6, 2010 U.S.
+Menkveld and Yueshen (2019) study the May 6, 2010 U.S.
 equity Flash Crash and argue that a disruption in cross-market arbitrage can make fragmented markets fragile and result in price crashes; they also refer to analogous flash-crash episodes in cryptocurrency markets, including in Ethereum (June 21, 2017) and Bitcoin (October 10, 2017).
 Because our baseline analysis is conducted at the weekly frequency using close-to-close returns, these very short-lived intraday dislocations are less likely to affect our main results, though they remain relevant for studies of market microstructure and tail risk.
-[^17] Barbon & Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
+[^16] Barbon and Ranaldo (2021) compare centralized and decentralized cryptocurrency exchanges along two key dimensions—market liquidity and price efficiency—and find that decentralized exchanges exhibit lower efficiency.
 Table 6, Panel A, reports the properties of an “arbitrage” (CARB) strategy that attempts to exploit these price discrepancies.
@@
 Not surprisingly, the mean excess returns are implausibly high—on the order of 68.5% per week over 2014–2025, rising to 95% per week from 2020 onward—precisely because the required arbitrage trades are blocked or prohibitively costly.
-Specifically, Borri & Shakhnov (2022) show that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe.
-That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer & Vishny (1997).
+Specifically, Borri and Shakhnov (2022) show that this seemingly exceptional performance is concentrated in the extreme portfolio that holds pairs with the highest Bitcoin price relative to the benchmark quote, where segmentation and frictions are most severe.
+That is, the cryptocurrency market inefficiencies reflect the classic market segmentation and trading frictions as in Shleifer and Vishny (1997).
 As Sophocles writes in Antigone, “Nothing vast enters the life of mortals without a curse.” Early cryptocurrency investors often told a different story: that outsized returns came from cross-exchange “arbitrage.” Public interviews with Sam Bankman-Fried—the founder of the now-defunct exchange FTX—describe how Alameda Research, a hedge fund he helped run, allegedly began by focusing on such trades.
 [^17] See, for example, Bloomberg’s coverage of the so-called Kimchi premium, which denotes the fact that it was more expensive to buy Bitcoin in Korean Won than in US dollars: https://www.bloomberg.com/news/articles/2024-11-15/south-korea-crypto-boom-what-is-the-kimchi-premium, and Lewis (2023)’s book about Sam Bankman-Fried.
-However, Borri & Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
-We revisit the analysis in Borri & Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish złoty, Swiss franc, British pound, and U.S.
+However, Borri and Shakhnov (2022) show that implementability is limited to a narrow set of highly liquid pairs on reputable cryptocurrency exchanges without binding capital controls.
+We revisit the analysis in Borri and Shakhnov (2022) and construct the CARB strategy for 226 Bitcoin pairs, traded on 96 exchanges against 11 fiat currencies (Australian dollar, Canadian dollar, Danish krone, euro, Hong Kong dollar, Israeli shekel, Japanese yen, Polish złoty, Swiss franc, British pound, and U.S.
 dollar).
@@
 These returns are not pure arbitrage—they are compensation for risk.
-Borri & Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
-The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu, Makarov & Schoar (2023), is the most visible manifestation of the results described in this fact.
+Borri and Shakhnov (2022) show that the expensive pairs (those priced above the benchmark) are systematically riskier—they depreciate more in bad times for crypto-investors, when aggregate liquidity and sentiment are low.
+The collapse of Alameda Research and the failure of FTX in November 2022—only a few months after the demise of the synthetic stablecoin ecosystem Terra–Luna described in Liu, Makarov, and Schoar (2023), is the most visible manifestation of the results described in this fact.
 More broadly, the evidence from cryptocurrency markets provides a clean illustration of the classic limits-to-arbitrage view: when trading, settlement, and capital-mobility frictions are economically large, violations of the law of one price can persist rather than being immediately arbitraged away.
@@
 Panel A uses the full sample of 510 BTC–fiat pairs traded on 165 centralized exchanges in 49 currencies.
-Panel B restricts to 226 BTC–fiat pairs on 96 exchanges against 11 fiat currencies, which are the most liquid fiat currencies on the most reputable exchanges in jurisdictions with minimal capital controls as in Borri & Shakhnov (2022).
+Panel B restricts to 226 BTC–fiat pairs on 96 exchanges against 11 fiat currencies, which are the most liquid fiat currencies on the most reputable exchanges in jurisdictions with minimal capital controls as in Borri and Shakhnov (2022).
 The post-2020 sample is the sample since Jan 1, 2020.
@@
 Stylized fact 6: Cryptocurrency carry trade, the defining characteristics of the futures market, used to be highly profitable, but profitability has compressed sharply since 2024.
-Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black & Scholes 1973; Merton 1973; Black 1976).
+Derivatives are a cornerstone of modern markets, enabling hedging, leverage, and price discovery across equities, interest rates, and commodities (Black and Scholes 1973; Merton 1973; Black 1976).
 There is widespread usage of derivatives in the cryptocurrency market, albeit with novel designs.
 Among all derivatives, perpetual futures are the most common derivative contracts in cryptocurrency markets.
-For example, Schmeling, Schrimpf & Todorov (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample.
+For example, Schmeling, Schrimpf, and Todorov (2023) document that perpetual futures account for more than 98% of the Bitcoin futures volume in their sample.
 Perpetuals, inspired by Shiller (1993), have no maturity; positions are marked to market (typically every 8 hours) and a funding rate—positive or negative by design—flows between longs and shorts to keep futures prices anchored to spot.
-Ackerer, Hugonnier & Jermann (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures.
-The crypto-carry trade strategy of Schmeling, Schrimpf & Todorov (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
+Ackerer, Hugonnier, and Jermann (2024) provides the theoretical foundations for the pricing of perpetual cryptocurrency futures.
+The crypto-carry trade strategy of Schmeling, Schrimpf, and Todorov (2023) is short a perpetual futures contract and long a position in the corresponding spot market.
 Figure 2 plots the evolution over time of one dollar invested in the crypto-carry strategy (black line, left y-axis) and in the spot Bitcoin market (red line, right y-axis).
@@
 The figure plots the cumulative returns from Tether Carry Trades (left axis) and from a long buy-and-hold strategy in Bitcoin (right axis).
-The Carry Trade is defined as in Schmeling, Schrimpf & Todorov (2023).
+The Carry Trade is defined as in Schmeling, Schrimpf, and Todorov (2023).
 The sample goes from August 1, 2020 to May 31, 2025.
@@
 When one party possesses superior information, such as inside information, prices may deviate from fundamentals, and capital may be misallocated (e.g., Akerlof 1970).
-Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983; Dye 1985; Healy & Palepu 2001; Verrecchia 2001).
-However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001; Ma & Shleifer 2025).
+Accounting research builds directly on this and shows how disclosure, regulation, and intermediaries can reduce these frictions (Verrecchia 1983; Dye 1985; Healy and Palepu 2001; Verrecchia 2001).
+However, in the absence of regulation and supervision, managers may strategically disclose for private benefits (Verrecchia 2001; Ma and Shleifer 2025).
 Policymakers thus issue regulations to address these inefficiencies and to produce credible, comparable, and decision-useful information.
-Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts & Zimmerman 1983).
+Intermediaries, such as analysts and auditors, emerge to reduce information asymmetry by analyzing and validating information (Watts and Zimmerman 1983).
 Information asymmetry remains pervasive in cryptocurrency markets, though regulation and disclosure have increased in recent years.
@@
 2025).
-On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George & Patatoukas 2021; Liu, Sheng & Wang 2025; Du, Huang & Wu 2025).
-Howell, Niessner & Yermack (2020) and Bourveau et al.
+On the cryptocurrency side, voluntary disclosures, such as whitepapers and social media announcements, have become increasingly common (e.g., George and Patatoukas 2021; Liu, Sheng, and Wang 2025; Du, Huang, and Wu 2025).
+Howell, Niessner, and Yermack (2020) and Bourveau et al.
 (2022) show the importance of voluntary disclosures in the initial coin offering stage.
 Studies also examine information intermediaries (e.g., analysts and auditors) and find that they play an important role in this market (e.g., Bourveau et al.
-2022; Bourveau, Brendel & Schoenfeld 2024).
+2022; Bourveau, Brendel, and Schoenfeld 2024).
 In July 2025, Congress enacted the Guiding and Establishing National Innovation for U.S.
@@
 households, showing that cryptocurrency investors differ systematically from other investors.
-Alvarez, Argente & Van Patten (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption.
-Han, Liu & Sui (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets.
-Finally, Benetton & Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies.
+Alvarez, Argente, and Van Patten (2023) provide complementary evidence on the use of cryptocurrencies as a means of payment, using the case of El Salvador—where Bitcoin was made legal tender—and show that privacy and security concerns remain key barriers to adoption.
+Han, Liu, and Sui (2023) study how social interactions and sentiments affect retail investor trading in cryptocurrency markets.
+Finally, Benetton and Compiani (2024) incorporate individual-level survey data into a structural characteristics-based demand model where investors can hold cryptocurrencies.
 They find that individual beliefs of cryptocurrency are different and are critical in explaining cryptocurrency valuation.
@@
 A significant segment of the cryptocurrency market, however, is non-fungible tokens (NFTs).
-Goetzmann & Nozari (2022) are the first to construct repeated sales return indices of NFTs and examine their properties.
+Goetzmann and Nozari (2022) are the first to construct repeated sales return indices of NFTs and examine their properties.
 Borri et al.
@@
 (2022a) show that, due to the high non-fungibility of NFTs, the market is highly segmented, which has implications for returns and investor behaviors in the market.
-Huang & Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing.
-Oh, Rosen & Zhang (2025) document preference-driven herding in NFT primary market sales.
+Huang and Goetzmann (2023) document that behavioral biases—such as selection neglect and extrapolative beliefs—play a significant role in NFT pricing.
+Oh, Rosen, and Zhang (2025) document preference-driven herding in NFT primary market sales.
 A stream of work examines the so-called dark side of cryptocurrencies, especially in the early years.
-Foley, Karlsen & Putniņš (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov & Schoar (2021) that show the results of Foley, Karlsen & Putniņš (2019) are largely driven by double-counting.
-Griffin & Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation.
-Amiram, Jørgensen & Rabetti (2022) show that abnormal spikes in on‑chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, Cong et al.
+Foley, Karlsen, and Putniņš (2019) argue that in the early days, the majority of Bitcoin transactions were about illegal activities, but this is recently challenged by Makarov and Schoar (2021) that show the results of Foley, Karlsen, and Putniņš (2019) are largely driven by double-counting.
+Griffin and Shams (2020) analyze blockchain and market data to argue that Tether flows timed after market downturns are associated with large Bitcoin price increases during the 2017 boom, consistent with price manipulation.
+Amiram, Jørgensen, and Rabetti (2022) show that abnormal spikes in on‑chain transfers around major terrorist attacks carry predictive power for attack timing and highlight chokepoints visible in public ledgers, Cong et al.
 (2025) provide an anatomy of crypto-enabled crimes and identify ransomware as the dominant activity, while Cong et al.
 (2023a) review crypto-related scams, including Ponzi schemes and “rug pulls” in decentralized projects and token offerings.
-Li, Shin & Wang (2025) document pervasive cryptocurrency pump‑and‑dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre‑event run‑ups consistent with insider gains at outsiders’ expense.
+Li, Shin, and Wang (2025) document pervasive cryptocurrency pump‑and‑dump episodes that generate brief surges in prices, volume, and volatility followed by rapid reversals, with pre‑event run‑ups consistent with insider gains at outsiders’ expense.
 5 CONCLUSIONS